### v1

# From CVE Entries to Verifiable Exploits: An Automated Multi-Agent Framework for Reproducing CVEs

Abstract
High-quality datasets of real-world vulnerabilities and their corresponding verifiable exploits are crucial resources in software security research. Yet such resources remain scarce, as their creation demands intensive manual effort and deep security expertise. In this paper, we present CVE-Genie, an automated, large language model (LLM)-based multi-agent framework designed to reproduce real-world vulnerabilities, provided in Common Vulnerabilities and Exposures (CVE) format, to enable creation of high-quality vulnerability datasets. Given a CVE entry as input, CVE-Genie gathers the relevant resources of the CVE, automatically reconstructs the vulnerable environment, and (re)produces a verifiable exploit. Our systematic evaluation highlights the efficiency and robustness of CVE-Genie’s design and successfully reproduces approximately 51% (428 of 841) CVEs published in 2024-2025, complete with their verifiable exploits, at an average cost of \$2.77 per CVE. Our pipeline offers a robust method to generate reproducible CVE benchmarks, valuable for diverse applications such as fuzzer evaluation, vulnerability patching, and assessing AI’s security capabilities.
1 Introduction
Tens of thousands of software vulnerabilities are discovered every year, with more than 49,000 vulnerabilities tracked by the National Vulnerability Database (NVD) [42] in 2024 alone. Many of these vulnerabilities are cataloged using the CVE format, which provides a standardized identification system to help organizations address security flaws in both software and hardware. Still, existing vulnerability detection tools, such as those listed by OWASP [44], struggle to keep up with the increasing volume of code released each year. As a result, many vulnerabilities remain undiscovered in codebases for years (e.g., up to 2 years in Chromium and 7 years in OpenSSL [3]) posing significant security risks.
A key requirement for developing effective automated vulnerability detection techniques is the availability of high-quality datasets with detailed examples and ground truth. Studies have shown that low-quality data can cause machine learning (ML) based vulnerability detection techniques to learn spurious correlations [4, 49]. Although public repositories, such as the NVD, include hundreds of thousands of CVEs, they often lack critical details, such as the vulnerable code, setup instructions for the vulnerable environment, working exploits, and validation steps, which are all required for the effective replication of a vulnerability and to build suitable ground truth datasets for research [38]. This is because the main purpose of vulnerability advisories is to alert system administrators about software that needs to be updated, and concrete exploit code is often kept confidential for ethical reasons. Unfortunately, generating this missing information from CVEs alone requires both substantial manual effort and extensive security expertise. For instance, Mu et al. [38] spent over 3,600 hours of manual effort with 43 security experts to reproduce publicly reported CVEs but managed to reproduce only 368 memory-related vulnerabilities in Linux.
Figure 1: CVE-GenieOverview.
To address these challenges, the security community has adopted several approaches: manual annotation [21, 65], inserting bugs into real-world code [41, 36], and heuristic-based mining [13, 10, 40, 15, 7]. While these methods have advanced the field, they come with limitations, such as limited support for programming languages and Common Weakness Enumerations (CWE) categories, mislabeled vulnerabilities [50], and data becoming stale. For instance, the DARPA Cyber Grand Challenge dataset from 2016 [11] quickly became outdated due to its small size and low-complexity bugs [20]. On the other hand, automated data collection helps address some of these limitations, but often tend to be limited in scope. For instance, ARVO [34] automatically curates reproducible builds with triggering inputs from OSS-Fuzz reports, but its scope is limited to memory corruption bugs in C/C++. This highlights a clear need for a dataset curation method that enables fully reproducible vulnerability instances across a wide range of programming languages and vulnerability types.
Recently, LLMs have demonstrated remarkable capabilities in solving complex software engineering (SWE) and security tasks, ranging from automated bug detection [39, 16] and repair [46, 57, 24], to writing fuzzer test harnesses [12]. Moreover, the shift to agentic workflows [63, 8, 62, 32], in which LLM-powered agents are provided access to real-world tools [1], has further pushed the capabilities of general purpose LLMs and lead to significant performance gains on extremely difficult real-world SWE related benchmarks, such as SWE-bench [25]. This raises the question whether LLM agents could also be useful in automatically setting up and reproducing existing CVEs, allowing us to build reliable, large, and realistic benchmarks for the security community.
In this paper, we present CVE-Genie, a fully-automated, LLM-driven multi-agent framework for end-to-end CVE reproduction, which embodies what we define as key properties of an ideal CVE reproduction system, captured by the acronym EAGER. Specifically, an ideal system: Generates working Exploit/proof-of-concept (PoC); Builds Assessors/verifiers for the exploit; Generalizes across CWEs, languages, and projects; Is End-to-end automated; and Rebuilds vulnerable environments. This paper makes the following contributions:
  1. 1.
    We develop CVE-Genie, a four-stage, end-to-end automated pipeline; Processor, Builder, Exploiter, and CTF Verifier, that extracts CVE data (i.e., source code, security advisories, patches, etc.), rebuilds the environment, generates exploits and verifiers to assess the exploits, and validates the final results.
  2. 2.
    We systematically select the optimal configuration of LLMs for CVE-Genie for optimal performance over diverse CVEs, encompassing diverse range of vulnerabilities, programming languages, and projects, including scenarios with incomplete CVE information.
  3. 3.
    We ran CVE-Genie on 841 CVEs (published between June 2024–May 2025), successfully reproducing 428 CVEs across 267 projects, 141 CWEs, and 22 programming languages, highlighting our generalizability.
  4. 4.
    We systematically evaluate CVE-Genie’s design and demonstrate that each component is essential for optimal performance: even advanced LLMs like o3 fail to reproduce even one CVE without our agentic framework.
  5. 5.
    We open-source our framework, source code, and datasets of reproduced CVEs as well as logs for all agents conversations, providing an ongoing valuable contribution to the community for downstream research tasks (e.g., tools for automated vulnerability detection).
2 Background and Related Work
Name Source
Real
# Projects
# CWEs
# Lang
Automated
Reproducible
PoC
Verifier
SARD [41] Manual
~120 ~150 6
SVEN [21] Manual
n/a 9 3
Devign [65] Manual
4 n/a 2
SecLLMHolmes [54] Manual
4 8 3
VulChecker [36] Manual
n/a 5 2
Draper [52] SA tool
n/a 9 2
D2A [64] SA tool
6 7 2
VUDENC [56] Sec. issue
812 7 1
DiverseVul [10] Sec. issue
797 150 12
PrimeVul [13] Patch diff
755 140 2
BigVul [15] Patch diff
348 91 2
CrossVul [40] Patch diff
1,675 168 40
CVEfixes [7] Patch diff
1,754 180 30
ARVO [34] OSS-Fuzz
273 4 2
CVE-Bench [66] Public CVEs
36 8 6
Mu et al. [38] Public CVEs
1 4 1
CVE-Genie Public CVEs
267 141 22
Table 1: Overview of vulnerability benchmark datasets. Real: synthetic , real-world , mixed . CVE-Genie results reflect CVEs from Jun 2024–May 2025.
Existing Benchmarks and Limitations. To evaluate various vulnerability detection techniques (rule-based methods [44, 58], ML-approaches [36, 30, 31, 29], and LLM-based systems [17, 19, 47, 18]), the community needs standardized benchmarks. Prior work has explored four main approaches for addressing these challenges. (1) Manual efforts involve analyzing [21] or reproducing [66] vulnerabilities in real-world code, or crafting synthetic examples [54], but datasets are small, lack ecological validity, and often suffer labeling errors due to human mistakes (e.g., 50% data in Devign [65] have wrong labels  [50]). (2) Automated labeling using static analysis (SA) tools, such as bandit [6] or infer [23], reduces manual effort but introduces many false positives [48, 64]. (3) Mining real-world data, such as GitHub issues or CVE patch commits, assumes that all removed or modified code is vulnerable; However, this assumption often leads to mislabeled samples, as Risse et al. [50] highlighted for DiverseVul [10] and BigVul [15]. (4) Relying on proven sources, such as OSS-Fuzz, offers verified exploits and sanitizer feedback; However, these are restricted to certain bug types, primarily memory corruption in C/C++, and do not generalize across languages or vulnerability classes [34].
Besides data quality, existing datasets also lack necessary artifacts for dynamic evaluation beyond static labels [4, 49, 54], including dynamic execution environments and reproducible exploits, which is difficult to obtain [38]. Table 1 provides an overview of the current state of vulnerability benchmarks.
Call for EAGER Style for CVE Reproduction. A CVE ID is a unique identifier for publicly disclosed security vulnerabilities in real-world software. Each CVE ID corresponds to a CVE entry that includes key details for that CVE, such as vulnerability description, source code, affected versions, security advisories, CWE classifications, and patches, enabling security professionals to assess and mitigate risks. Creating a dataset from the standard CVE database is a promising way towards addressing the aforementioned challenges (see Appendix B). Therefore, we introduce EAGER, a set of crucial criteria for ideal CVE reproduction:
  • Exploit Generation – (re)creation of an exploit or PoC that reliably triggers the vulnerability in the CVE.
  • Assessment – inclusion of a “verifier” or “sanitizer” capable of assessing whether the generated exploit or PoC successfully triggers the vulnerability.
  • Generalization – reproduction of CVEs across diverse CWEs, programming languages, and software projects.
  • End-to-end Automation – execution of all stages of CVE reproduction in a fully automated manner.
  • Rebuild project – reconstruction of the original vulnerable environment to facilitate exploit/PoC execution.
However, none of existing benchmarks satisfy all the criteria (Table 2). For example, manual effort for producing EAGER datasets does not scale and is not generalizable [38, 66]. ARVO [34] focuses solely on memory-related bugs from OSS-Fuzz C/C++ projects, limiting its scope. In contrast, CVE-Genie is the only framework satisfying all criteria.
Method Exploit Assess General E2E Rebuild
Manual [32, 66]
Fuzzers
Metasploit [35]
ARVO [34]
CVE-Genie
Table 2: Comparing Potential Vulnerability Reproduction Methods and their EAGER attributes.
Large Language Models and Agents. LLMs are transformer-based neural network models with billions of parameters. They have shown remarkable performances in various application domains, such as question answering [43], mathematical reasoning [2], and code generation [9, 5]. Recent work also shows their utility in security tasks, e.g., such as bug detection and repair [46, 57, 24] and writing fuzzer test harnesses [12].
LLM agents are systems that combine LLMs with tools. They can finish complex tasks via a sequence of actions, including planning, generation, and tool executions [53, 22, 45, 27, 26]. An agentic system can have multiple agents, each with a sub-task, such as reasoning, critiquing, generating, perceiving, remembering, or teaching, as well as its own workflow and tool sets. In security, researchers have started to construct LLM agents to find and repair vulnerabilities. There are multi-agent patching systems, where different sub-agents are responsible for fault localization, patch generation and validation (e.g., PatchPilot [28], and PatchAgent [61]). Through the recent DAPRA AIxCC competition, researchers constructed agentic-based systems with end-to-end capabilities for vulnerability detection and patching [18]. These recent successes provides the argument for using LLM agents in vulnerability analysis, software development, and external tool interaction, key capabilities for building an EAGER-style automated end-to-end CVE reproduction framework.
3 CVE-Genie
Like human analysts [38], CVE-Genie utilizes four core modules to process and reproduce CVEs: (1) the Processor (Section 3.1), which retrieves source code and constructs a structured knowledge base; (2) the Builder (Section 3.2), which reconstructs the vulnerable environment; (3) the Exploiter (Section 3.3), which reproduces the exploit; and (4) the CTF Verifier (Section 3.4), which generates a verifier for the exploit. We also design and integrate a set of tools that enable LLM agents in these modules to perform their respective tasks (see Appendix A). To illustrate the framework in practice, we also present a full end-to-end reproduction of CVE-2024-4340 by CVE-Genie (Figure 2).
To bridge the gap identified in Section 2, namely the lack of any system capable of true EAGER-style CVE reproduction, CVE-Genie’s design is guided by the following principles.
  1. 1.
    Modular Task Decomposition: Since LLMs struggle with long, complex contexts [54] and tasks [25], CVE-Genie decomposes reproduction into focused modules handled by specialized agents with systematically engineered prompts [^1] All prompts will be provided in the open-source artifact..
  2. 2.
    Robustness to Incomplete Data: Given that missing CVE fields can reduce reproduction success by up to 44% [38], agents in CVE-Genie can fall back on patch analysis and direct source-code inspection when advisories or PoCs are unavailable.
  3. 3.
    Reliability through Self-Critique: To counter LLMs’ reasoning limits [54], each module uses paired developer and critic agents in a ReAct-style loop [60], enabling iterative refinement and internal verification.
Together, these principles ensure CVE-Genie can deliver the first practical, automated, and generalizable framework for EAGER-style CVE reproduction. The following subsections detail the workings of each CVE-Genie component, using CVE-2024-4340 as a running example to illustrate the end-to-end reproduction shown in Figure 2.
Figure 2: CVE-Genie architecture and an end-to-end example of workflow of reproduction for CVE-2024-4340, i.e., Denial of Service due to RecursionError in sqlparse < v0.5.0.
3.1 Processor
This module comprises of two sub-modules: the Data Processor, which locates the vulnerable project’s source code and gathers relevant raw resources, and the LLM-based Knowledge Builder, which transforms these resources into a structured knowledge base. Below we provide detailed functionalities of these sub-modules.
3.1.1 Data Processor 1
Goal/Rationale. For the given CVE ID, this sub-module identifies and extracts the vulnerable version of the software along with all relevant CVE details, as follows:
Source Code Extraction. The Data Processor begins by locating the source code for the vulnerable project linked to the CVE, an essential requirement for CVE-Genie to reproduce the vulnerability. It searches the GitHub URLs referenced in the “cvelist”222cvelist” is an automated pilot program for CVE submission through GitHub (https://github.com/CVEProject/cvelist) to identify the project repository. For closed-source software, Data Processor provides the option to the user to supply the repository URL. Once the source code is located, it identifies the affected software configurations: specific versions, platforms, or settings vulnerable to the CVE. For instance, CVE-2024-4340 affects all versions of “sqlparse” prior to v0.5.0. Based on this, the latest affected (vulnerable) version (v0.4.4) is retrieved using the GitHub API, and its source code is downloaded (as shown in Figure 2).
Vulnerability Information Extraction. After the vulnerable version of the project source code is downloaded, Data Processor extracts the following four key pieces of information from “cvelist”, if available (see Figure 2):
  1. 1.
    CVE description: A high-level summary of the vulnerability in the target project.
  2. 2.
    CWE data: The CWE information, a categorization system for hardware and software weaknesses that associates the given CVE with a speficic type of vulnerability, e.g., CWE-674 (Uncontrolled Recursion) [^3] https://cwe.mitre.org/data/definitions/674.html.
  3. 3.
    Patch commits: The git-diff of code changes in the target source code that fix the vulnerability. These commits help LLMs localize the issue, understand its root cause [21, 13, 15], and generate effective exploits.
  4. 4.
    Security advisories and PoC: We filter URLs referenced in the CVE entry by keywords (e.g., “security”, “advisory”, “advisories”, “bounty”, “bounties”), and scrape their contents.
Output. The vulnerable version of a project’s source code alongside its directory structure and CVE raw data.
3.1.2 Knowledge Builder 2
Goal/Rationale. This agent is instructed to analyze, distill, and organize the raw information of, extracted by the Data Processor, into a structured knowledge base while retaining essential details required to reproduce the given CVE. This includes extracting PoC details or step-by-step instructions from security advisories, which are invaluable for accurately recreating the vulnerability. For instance, CVE-2024-4340 has a PoC provided in its security advisory and the patch commit helps understand the root cause by highlighting where the vulnerability was mitigated. As shown in Figure 2, the Knowledge Builder included both of these details in the knowledge base. In case an advisory does not include an exploit, the LLM agent is tasked with generating an overview of what an exploit might entail. Additionally, capturing the patch details helps to localize the root cause of the vulnerability [21, 13, 7] and aids in crafting an effective exploit. This knowledge base acts as the long-term memory for agents, helping them to effectively reproduce CVEs. By including only essential information, the knowledge base avoids overloading the LLM’s context window, ensuring efficient reproduction.
Output. A structured CVE knowledge base providing details of CVE including its associated CWEs, affected software configurations, root causes (if provided in the security advisory), and essential information from security advisories and patch commits necessary for exploitation.
3.2 Builder
Once the knowledge base is populated and the vulnerable version of the project’s source code is downloaded, the next objective is to build the project in a way that allows the exploit to be executed. To achieve this, our pipeline employs two developer LLM agents: the Pre-requisite Developer Agent, which analyzes and plans the environment setup, and the Setup Developer Agent, which executes this plan to configure the vulnerable environment. As well as one critic, the Setup Critic Agent, which analyzes the logs of the Setup Developer and evaluates the setup for the project, allowing the system to correct its mistakes. Below, we discuss the functionality of these agents.
3.2.1 Pre-Requisite Developer Agent 3
Tools. For this step, we only provide the LLM agent with access to read-only tools, i.e., execute_ls_command, and get_file_by_name, which the agent can use to traverse the code base and read files, because we do not want the agent to execute any commands or write to files in this phase.
Goal/Rationale. We introduce this agent as an initial exploration and planning step before actually setting up the vulnerable project. This is done for three key reasons.
(1) Our preliminary experiments reveal that many projects contain inaccurate and incomplete information in their README files (e.g., vulnerable version v1.2.7 of lunary for CVE-2024-5129, includes incorrect paths to .env files that are crucial for the setup). Without the Pre-Requisite agent, these details were consistently overlooked by LLMs, causing setup failures. Identifying and correcting such issues early reduces unsuccessful attempts, conserves the LLM’s context window, and improves the likelihood of successful environment setup.
(2) During environment setup, the agent’s job is to explore and analyze the source code to identify critical components and to build the project by executing a series of commands. For large projects, having to do both tasks with one agent can exceed the context limit and reduce efficiency (as demonstrated in the ablation study in Section 4.5). To address this, we delegate the exploration of code base to a dedicated agent, the Pre-Requisite Developer, allowing the setup process to focus more on building the project.
(3) The Pre-Requisite Developer also defines the “expected state,” i.e., when the project is fully set up and ready for exploitation. This state guides the Setup Developer Agent in verifying the correct configuration of the vulnerable environment. For example, in CVE-2024-4340 (Figure 2), the Pre-Requisite Developer explored the root directory and key files such as README and sql.py to understand the project context, then specified the expected state as having sqlparse version v0.4.4 installed for the exploit to work correctly.
Output. (1) Detailed overview of the project, (2) important files to pay attention to during the setup, (3) required services and their configurations, and (4) expected state of the project upon successful setup to facilitate verification.
3.2.2 Setup Developer Agent 4
Tools. For this step, we allow access to the following tools, i.e., get_file, write_to_file, execute_ls_command, execute_linux_command, and set_environment_variable.
Goal/Rationale. The Setup Developer begins in the directory containing the vulnerable version’s source code and focuses on configuring the project based on instructions from the Pre-Requisite Developer, exploring the codebase when necessary. For CVEs involving third-party libraries, our initial experiments showed that setup is greatly simplified by using package managers like pip or npm to install specific vulnerable versions instead of building from source. For example, in the case of CVE-2024-4340 affecting sqlparse, the Pre-Requisite Developer instructed the installation of version 0.4.4, and the Setup Developer executed pip install sqlparse==0.4.4 directly (command 4a, Figure 2), resulting in a smooth setup. If the package manager fails, the Setup Developer falls back to building from source. Typically, the Pre-Requisite Developer also recommends running a basic PoC to confirm readiness before handing off to the next agent (e.g., commands 4e and 4f in Figure 2).
Output. This agent returns a final decision whether the setup is successful or not. If successful, the output also includes instructions on how another agent can access the running project.
3.2.3 Setup Critic Agent 5
Goal/Rationale. The Setup Critic evaluates whether the project setup performed by the Setup Developer is correct and complete, ensuring that the environment is properly configured for the vulnerability described in the CVE knowledge base to be exploited. Additionally, this agent plays a critical role in detecting deceptive or shortcut behaviors by the LLM, such as fabricating a mock project instead of performing a genuine setup, or deliberately injecting vulnerabilities into the codebase to simplify exploitation. These undesirable patterns are discussed in detail in Section 4.1.1, which informs the instruction design for the Setup Critic. Based on this, the critic is equipped to identify both functional and security-related flaws in the setup process.
Output. A comprehensive analysis of the setup logs, a binary decision on whether the setup is valid and complete, and actionable feedback for correcting issues or improving the setup if necessary.
1
2"""
3PoC CVE-2024-4340 (sqlparse < 0.5.0)
[…]
5python3 exploit.py 10000
6"""
7import sys
8import sqlparse
[…]
10if len(sys.argv) != 2:
11print("Usage: python3 exploit.py <depth>")
12sys.exit(1)
13arg = sys.argv[1]
14try:
15depth = int(arg)
16payload = "[" * depth + "]" * depth
17except ValueError:
18payload = arg
19sqlparse.parse(payload)
[…]
21main()
(a) Exploit script for CVE-2024-4340. It constructs a deeply nested list payload and submits it to sqlparse.parse(), triggering uncontrolled recursion. The included comment documents a crashing input (depth 10,000) that reliably triggers the vulnerability.
1
2Traceback (most recent call last):
3File "poc.py", line 20, in <module>
4sqlparse.parse(payload)
5....
6File "sqlparse/sql.py", line 214, in flatten
7yield from token.flatten()
8[Previous line repeated 983 more times]
9RecursionError: maximum recursion depth exceeded
(b) Execution traceback produced by the exploit, showing a RecursionError after excessive recursive calls within sqlparse.
Figure 3: PoC for CVE-2024-4340 in sqlparse v0.4.4, showing the exploit script and the resulting crash.
3.3 Exploiter
After successfully configuring the vulnerable system, the next step in our pipeline is to generate an exploit for the given vulnerability. CVE-Genie accomplishes this using two LLM agents: the Exploit Developer, which develops the PoC for an exploit in the vulnerable environment, and the Exploit Critic, which analyzes the logs of the Exploit Developer and evaluates the exploit for the given vulnerability. Below, we discuss the detailed functionality of these two agents.
3.3.1 Exploit Developer Agent 6
Tools. This agent uses the same tools as the Setup Developer.
Goal/Rationale. The Exploit Developer is responsible for crafting and demonstrating exploits for vulnerabilities listed in the CVE knowledge base on pre-configured vulnerable systems. If a PoC is provided, the agent replicates and verifies it by executing the script with the appropriate triggering input before delivering the final, working PoC. For example, in CVE-2024-4340 (Figure 2), the PoC was available, so the agent validated the setup (6b), demonstrated the exploit (6c) by triggering RecursionError in sqlparse/sql.py (see Figure 3(b)), and submitted the verified PoC script (see Figure 3(a)). If no PoC or exploitation steps are available, the Exploit Developer iteratively analyzes the codebase using available tools to develop a functional exploit. Upon success, it produces a Python PoC script that accepts the crashing input via command-line, includes comments on expected input format, and provides an example input. This ensures that the CTF Verifier can reliably validate the exploit without repeating the analysis.
Output. The agent returns its final decision whether it considers the exploit is successful or not. If successful, it provides an overview of the exploit and a Python PoC script.
3.3.2 Exploit Critic Agent 7
Similar to the role of the Setup Critic, the Exploit Critic is tasked with evaluating the behavior of the Exploit Developer by analyzing its execution logs. Its primary goal is to determine whether the generated exploit is both valid and of high quality. This includes verifying that the exploit could plausibly succeed in a real-world setting; that it does not rely on manipulating or altering the original setup in illegitimate ways; and that it avoids fabricated shortcuts, oversimplified logic, or artificial assumptions that would not hold in practice.
To perform this assessment, the Exploit Critic leverages a predefined set of behavioral patterns and common failure cases, as detailed in Section 4.1.2. By comparing the exploit’s behavior against these known issues, the critic can detect actions that are unfaithful to the original intent or technically incorrect. When flaws are found, the Exploit Critic provides structured, actionable feedback designed to help improve the robustness and realism of the exploit in subsequent iterations.
3.4 CTF Verifier
After generating a working exploit for a given CVE, the final step is to create a verifier script that confirms whether the exploit was successful. This approach is inspired by security Capture-The-Flag (CTF) challenges, where exploiting a vulnerability reveals a hidden “flag” i.e., a string of random characters or numbers. The CTF Verifier uses three sub-modules to develop and validate such verifiers. The Verifier Developer agent module generates a candidate verifier, the Flag Checker module ensures that the exploit correctly retrieves a flag, and the Verifier Critic agent assesses the quality and effectiveness of the verifier for the given exploit.
1
2import os, subprocess, sys
3FLAG = "3xploit66full"
[…]
5# Configuring PYTHONPATH
[…]
7cmd = [sys.executable, "exploit.py", "10000"]
8return subprocess.run(cmd, timeout=20)
[…]
10output = (proc.stdout or "") + (proc.stderr or "")
11return proc.returncode != 0 and "RecursionError" in output
[…]
13pre_setup()
14try: result = run_exploit()
15except subprocess.TimeoutExpired: sys.exit(1)
16if post_exploit(result):
17print(FLAG)
(a) Attempt 1 – Weak verifier script: pre_setup() lacks a check for the vulnerable sqlparse version. The exploit is run with a 20s timeout, risking missed RecursionErrors if they occur after the limit. The verification relies solely on detecting the string RecursionError in output, which can be trivially spoofed.
1
2import os, sys, importlib, traceback
3FLAG = "3xploit66full"
[…]
5assert sqlparse.__version__ < "0.5.0"
[…]
7sys.argv = ["exploit.py", "10000"]
8try:
9importlib.import_module("exploit").main()
10return "no_exception", ""
11except RecursionError: return "recursion_error", traceback.format_exc()
12except: return "other_exception", traceback.format_exc()
[…]
14return status == "recursion_error" and "sqlparse/sql.py" in tb
[…]
16pre_setup()
17status, tb = run_exploit_inprocess()
18if exploit_succeeded(status, tb):
19print(FLAG)
(b) Attempt 2 with Verifier Critic’s feedback – Improved verifier: Validates project version sqlparse < 0.5.0, runs the exploit in-process to avoid timeout issues, and explicitly checks for a genuine RecursionError with traceback originating from sqlparse/sql.py, preventing spoofed or misleading outputs.
Figure 4: Verifier scripts for exploit (in Figure 3) for CVE-2024-4340 corresponding to the run in Figure 2, illustrating the progression from a weak to a robust verifier based on Verifier Critic feedback.
3.4.1 Verifier Developer Agent 8
Tools. This agent follows the same read-only tools as the Pre-Requisite Developer agent, as in this step we don’t want the model to execute anything, and we do all validations using the Flag Checker and the Verifier Critic.
Goal/Rationale. Following the CTF methodology, we prompt the Verifier Developer to create a general CTF-style Python verifier script for a given PoC script. So, if the verifier script runs the PoC and it successfully triggers the vulnerability, it returns this flag (3xploit66ful). We prompt the Verifier Developer to format the verifier script in three structured steps:
  1. 1.
    Pre-Setup: Prepares the necessary inputs for the exploit, sets up the environment, and prepares a flag to return if the exploit correctly triggers the vulnerability.
  2. 2.
    Exploit Execution: Runs the provided PoC script to trigger the vulnerability. We make sure that the agent cannot modify the PoC script to avoid contamination.
  3. 3.
    Post-Setup: Verifies the success of the exploit and if successful, the script returns the flag.
This structured format ensures consistent and reliable validation across a broad spectrum of vulnerability exploit types.
Output. A Python script that acts as a custom verifier for the exploit/PoC of the given CVE.
3.4.2 Flag Checker 9
Once the verifier script is generated for the given exploit, before passing it to the critic agent, the Flag Checker executes the verifier script to ensure that it runs with the exploit script without any errors and produces the expected output i.e., the correct flag. For instance, figure 2 shows the correct flag 3xploit66ful is released when executing the verifier scripts generated for CVE-2024-4340. If a script fails due to a runtime error, or the flag is missing from the output, the system assumes that the verifier is incorrect. Feedback is then sent to the Verifier Developer to revise the script. This process is repeated iteratively, with a maximum of five attempts allowed to achieve functional correctness of the verifier script.
3.4.3 Verifier Critic Agent 10
Once the verifier and exploit scripts pass functional validation via the Flag Checker, they are further examined by the Verifier Critic, which performs a similar evaluative role as the other critic agents, focusing specifically on the response produced by the Verifier Developer. It examines whether the verification process was properly carried out, without assumptions, omissions, or unreliable heuristics. Using the behavioral issues identified in Section 4.1.3, the agent flags incomplete or misleading verification strategies and provides feedback on how the verification could be made more robust or accurate. For instance, Figure 4 shows a detailed example of verifiers generated for CVE-2024-4340 with a weak and bypassable verification criteria, which was corrected using the Verifier Critic’s feedback.
3.4.4 Store Reproduced CVE 11
If the given CVE passes the Verifier Critic check, CVE-Genie marks it as reproduced and stores the VM snapshot for the vulnerable environment as well as the exploit and verifier scripts. Moreover, CVE-Genie also stores the metadata such as LLM cost, time spent, and all agents conversations.
4 Experimental Methodology
We begin by identifying the most suitable LLMs for integration into CVE-Genie by evaluating ten leading models (Table 3) across CVE-Genie’s three core modules, Builder, Exploiter, and CTF Verifier, using a diverse set of 60 post-knowledge-cutoff CVEs (Section 4.1). Based on this evaluation, we select the optimal LLM configuration and conduct a four-part evaluation of CVE-Genie: (1) a baseline study assessing consistency and efficiency across various vulnerabilities, languages, and projects (Section 4.2); (2) robustness testing under incomplete CVE information (Section 4.3); (3) large-scale evaluation on 841 CVEs (Section 4.4); and (4) an ablation study to assess the impact of key components of CVE-Genie on its scalability and effectiveness (Section 4.5). These evaluations are guided by the following research questions:
  • RQ1: How does the performance of CVE-Genie vary with the types of vulnerabilities, languages, and projects?
  • RQ2: Can CVE-Genie effectively reproduce CVEs with limited information, and which CVE report components are most crucial for reproduction?
  • RQ3: Can CVE-Genie reproduce a broad range of CVEs across diverse CWEs, languages, and projects?
  • RQ4: Is the complex architecture of CVE-Genie necessary, or can standalone LLMs achieve comparable results?
Model # Params Max Input Max Output Reas-oning Open Src Cutoff
o3 n/a 200k 100k 05/2024
o4-mini n/a 200k 100k 05/2024
Claude 3.7 Sonnet n/a 200k 64k 11/2024
Claude 3.5 Sonnet 175B 200k 8k 04/2024
Gemini 2.5 Pro n/a 1M 65k 01/2025
Gemini 2.5 Flash n/a 1M 65k 01/2025
Llama 4 Maverick 400B 1M n/a 08/2024
Qwen 3 235B 128k 64k 06/2024
Deepseek V3 671B 64k 8k 07/2024
DeepSeek R1 671B 128k 8k 07/2024
Table 3: Studied LLMs.
4.1 Selecting Optimal LLMs for CVE-Genie
Dataset $(D_{S})$. To evaluate the capabilities of LLMs in reproducing vulnerabilities using CVE-Genie, we construct a small-scale, diverse dataset $(D_{S})$ consisting of $60$ CVEs. These CVEs were published after the latest knowledge cutoff date of the LLMs under study (January 2025), as listed in Table 3. To ensure that $D_{S}$ is both representative and diverse, we select 40 CVEs from the top 25 most dangerous CWE [37] categories, and the remaining 20 CVEs are randomly sampled from the other CWE types. The resulting dataset, $D_{S}$, comprises 60 CVEs drawn from 44 different CWE categories, spanning 18 programming languages, and involving 56 distinct software projects.
Methodology. CVE-Genie is composed of four modules. The Processor’s LLM-based component, the Knowledge Builder, only performs a trivial task of summarization, so we assign a lightweight LLM (o4-mini) to it. In contrast, the remaining modules i.e., Builder, Exploiter, and CTF Verifier, require strong security vulnerability reasoning and SWE capabilities. Thus for each of these modules, we systematically select the best performing LLM for “developer” and “critic” tasks using the following 5-step procedure:
  1. 1.
    Candidate Selection: For each module, we identify subsets of LLMs from Table 3 as potential developers and critics, based on prior evidence of their capabilities, as well as time and cost constraints.
  2. 2.
    Developer Evaluation: Each candidate developer LLM is run for the given module, and its outputs are manually scored by the authors. The highest-scoring LLM is chosen as the developer.
  3. 3.
    Critic Prompt Design: We analyze common mistakes made by developers and use these insights to design a task-specific critic prompt for the module, enabling the critic to effectively detect those mistakes.
  4. 4.
    Critic Evaluation: Each candidate critic LLM is evaluated on recall or true positive rate (TPR) and specificity or true negative rate (TNR). When TPR is equal, we prefer the model with higher TNR, and vice versa. The critic with the best trade-off is selected.
  5. 5.
    Feedback Loop Assessment: We rerun the chosen best developer–critic pair for the given module, and evaluate the correctness of critic’s feedback and the developer’s capability to correctly address the critic’s feedback.
4.1.1 Builder
(a) Developer Agent
(b) Critic Agent
Figure 5: Builder Optimal LLM Evaluation
Candidate Selection. Setting up a project from scratch is the most time-consuming and complex part of CVE-Genie, as security advisories or CVEs do not provide instructions for setting up the vulnerable project. To address this, we use two developer agents: one for exploring the project and another for setting it up (see Section 3.2). This process requires extensive tool calls and command executions, making it both slow and costly with reasoning-heavy models like o3 or claude-3.7-sonnet, which can cost \$10–20 per CVE. Therefore, we select fast, cost-effective LLMs with sufficient context capacity, o4-mini, claude-3.5-sonnet, gemini-2.5-flash, llama-4-maverick, deepseek-v3, and qwen3, as developer agent candidates. For critic agents, we opt for stronger reasoning models such as o3, o4-mini, claude-3.7-sonnet, gemini-2.5-pro, and deepseek-r1, as the critic’s task only requires one call, and therefore, they incur minimal overhead.
Developer Evaluation. We ran CVE-Genie up to the Setup Developer stage 4 on CVEs from $D_{S}$ and manually evaluated each developer LLM using the following procedure. (1) For each CVE, we first identified the official installation documentation for the software’s vulnerable version and, based on it, established both the expected final build state and the sequence of commands required for installation. (2) We then reviewed the logs of runs where the LLM reported successful setup, comparing the executed commands against the documented installation process to assess alignment and completeness. (3) Finally, we validated the resulting environment: for libraries or binaries, we confirmed the vulnerable version by invoking the corresponding --version (or equivalent) command (e.g., for CVE-2025-1215 we verified that vim v9.1.1096 was installed), while for server-based software we performed a health check to ensure the service was running and accessible at the expected port.
As shown in Figure 5(a), o4-mini emerged as the best developer LLM for the Builder, successfully setting up 32 out of 60 CVEs projects, while making an average of 20 tool calls per CVE. In contrast, claude-3.5-sonnet mostly gave up on the given setup task, and models like gemini-2.5-flash and open-source LLMs were unreliable: Qwen 3 often described setup steps without executing them using tool calls, while gemini-2.5-flash issued tool calls with frequent syntax errors, eventually leading to the failed setup of the given project.
Critic Prompt Design. From analyzing developer LLM setups, we observed three recurring mistakes: (1) when setup fails, the LLM builds a simplified mock-up substitute project instead; (2) for pip/npm packages, it installs the latest version instead of the specified vulnerable one; and (3) when a server is required, it assumes commands like npm run dev succeed without verifying server health. We designed the critic prompt to detect these errors while reviewing Setup Developer logs.
Critic Evaluation. As shown in Figure 5(b), the o3 model achieves the highest TPR and TNR when evaluating Setup Developer’s execution logs. Most of its false negatives involved setups requiring external hardware, which Exploit Developer cannot exploit anyway. Some valid setups were also mistakenly rejected due to limited post-setup checks. Given its strong balance of recall and precision, we choose o3 as the critic model for the Builder.
Feedback Assessment. We evaluated CVE-Genie up to Setup Developer stage 4, using the best-performing developer (o4-mini) and critic (o3) models. Minor issues flagged by the critic (e.g., limited verification) were typically resolved by the developer in one iteration. However, fundamental issues (e.g., mock-up version of the project) persisted even after five iterations. Therefore, we limit feedback to a single iteration to efficiently address minor issues without incurring excessive overhead.
4.1.2 Exploiter
(a) Developer Agent
(b) Critic Agent
Figure 6: Exploiter Optimal LLM Evaluation
Candidate Selection. Generating a working exploit from a given CVE in a pre-configured vulnerable environment is a crucial task for reproducing CVEs. However, the Exploiter requires less code base exploration than the Builder, as advisories often provide PoCs, exploitation steps, or vulnerability details. Thus we include expensive reasoning models like o3, claude-3.7-sonnet, and gemini-2.5-pro to the candidates list of developers. To mitigate instability seen in open-source models (Section 4.1.1), we employ only closed-source LLMs (Table 3) as developers, while retaining the same critic LLMs as in the Builder.
Developer Evaluation. For the 32 successful setups from Section 4.1.1, we ran CVE-Genie up to the Exploit Developer stage 6 and manually evaluated each developer LLM as follows: (1) we checked whether the LLM executed a PoC or command sequence that demonstrated the exploit, (2) we verified that the vulnerability was clearly triggered (e.g., uncontrolled recursion in CWE-674 resulting in a Maximum Recursion Error), (3) we assessed fidelity to the CVE description, (4) we ensured no mock-up environments or fake exploits were used, and (5) we confirmed the exploit script followed the expected format (see Section 3.3.1).
Among all models, o3 consistently performed best, successfully reproducing the highest number of exploits (13 out of 27) and demonstrating them end-to-end within the actual vulnerable environment. In contrast, o4-mini often produced incorrect or mock-up exploits, and struggled with runtime errors. While claude and gemini models showed strong capabilities in demonstrating exploits, they lacked versatility across different vulnerability types. Hence, we select o3 as the most effective Exploit Developer.
Critic Prompt Design. We examined the failed exploit attempts and categorized the recurring error patterns. These error types directly mirrored our manual verification criteria (e.g., failure to demonstrate the exploit, unclear triggering of the vulnerability, deviation from the CVE description, reliance on dummy/fake exploits, or incorrect script formatting). We then incorporated these categories into the critic agent’s prompt, allowing it to automatically detect the same issues in the Exploit Developer logs.
Critic Evaluation. As shown in Figure 6(b), o4-mini achieved the best balance of recall and specificity. In contrast, o3 was overly strict, rejecting correct exploits and focusing too much on code formatting, while gemini-2.5-pro frequently overlooked incomplete exploit verification. Claude-3.7-sonnet was too lenient, accepting nearly all exploits. Based on this evaluation, we selected o4-mini as the most effective critic model for the Exploit Critic.
Feedback Assessment. We evaluated feedback effectiveness by running CVE-Genie on 32 correct setups with the top developer (o3) and critic (o4-mini). This yielded valid exploits for 16 CVEs, of which 3 were fixed based on the critic’s feedback. As in Section 4.1.1, the developer mainly resolved minor issues (e.g., incomplete verification or missing log triggers), while more complex flaws (e.g., incorrect exploit verification) persisted. To balance usefulness and cost, we therefore restrict feedback to a single iteration, focusing on quick corrections of minor errors.
4.1.3 Verifier
(a) Developer Agent
(b) Critic Agent
Figure 7: Verifier Optimal LLM Evaluation
Candidate Selection. Like exploit generation, creating a verifier also demands deep understanding of security vulnerabilities to ensure the exploit is successfully triggered. Therefore, we use the same developer and critic candidates as in the Exploiter.
Developer Evaluation. For the 16 valid exploits from Section 4.1.2, we ran CVE-Genie up to the Flag Checker stage 9 and manually evaluated each developer LLM on three criteria: (1) whether the verifier script followed the required format (see Section 3.4.1), (2) whether it executed the provided exploit script without modification, and (3) whether its verification logic was precise and reliable (see Figure 4 for a detailed example).
As shown in Figure 7(a), the o3 model generated 13 valid verifiers out of 14, outperforming all others. While o4-mini achieved comparable numbers, its verifiers more often relied on weak logic. Other models struggled with functional correctness and error recovery. We therefore selected o3 for the Verifier Developer due to its consistent one-shot success in producing correct verifiers.
Critic Prompt Design. We analyzed failures in verifier generation and found that the recurring error types matched our verification criteria (e.g., incomplete checks, incorrect validation logic, or misuse of the target environment). We used these categories to design the critic agent’s prompt, enabling it to automatically flag such errors when evaluating verifiers produced by the Verifier Developer.
Critic Evaluation. As shown in Figure 7(b), only the o3 model successfully identified critical but subtle issues in the verifier scripts, such as weak or flawed verification logic (Figure 4). Other models provided weak or no critique, making o3 the most effective critic for the CTF Verifier.
Feedback Assessment. We evaluated CVE-Genie on 16 correct exploits from the Exploiter, successfully generating accurate verifiers for 15 of them, of which 4 verifiers were improved through critic feedback. In our experiment, the maximum number of feedback iterations required to resolve either functional or critique-based issues was 4; therefore, we allow up to 5 retries for both the Flag Checker and the Verifier Critic.
4.1.4 Time and Cost Constraints
In our evaluation of CVE-Genie, we found that successfully reproduced CVEs typically required about \$2 and 18 minutes on average, with the most resource-intensive case costing \$6 and taking 45 minutes. Failed reproductions, on the other hand, averaged up to \$4 and 35 minutes. To balance flexibility with cost control, we set a per-CVE budget cap of \$5 and a maximum runtime of 45 minutes for all experiments.
4.2 Diversity-Aware Baseline Assessment
Figure 8: CVE-Genie performance over five runs on $D_{S}$.
Dataset. For this study we use the same dataset $D_{S}$ of 60 diverse CVEs (see Section 4.1).
Methodology. Due to the non-deterministic nature of LLMs, a single run may fail due to incorrect reasoning, while later attempts might succeed [33]. To account for this, we run CVE-Genie multiple times on the CVEs in $D_{S}$, using its most optimal configuration (Section 4.1). After each iteration, we store successfully reproduced CVEs and rerun on the remaining ones. This iterative approach helps maximize CVE reproduction and reveals CVE-Genie’s convergence over seven runs.
Results. We use statistical methods as well as manual analysis of randomly selected 25 CVE reproduction runs, and we categories our results and analysis in the following categories:
Variance Over Multiple Runs. The variability in CVE-Genie’s performance primarily stems from the Builder phase, which is more open-ended than the targeted exploit generation guided by CVE context. Project setup involves repository exploration (Pre-Requisite Developer) and command execution (Setup Developer), where environment-specific behaviors often cause failures. For example, one run failed because the Setup Developer modified the PATH variable, while a clean retry succeeded, highlighting the value of reattempting in a fresh environment. Since the CTF Verifier depends on the Exploiter, which in turn relies on the Builder, this initial project setup is a critical bottleneck. As a previously failed build, if later successful, can enable downstream modules to reproduce a new CVE (as illustrated in Figure 8). Over multiple runs, the number of successful project, exploit, and verifier builds showed consistent gains with convergance at run seven, indicating that most recoverable failures can be resolved within a few retries.
Vulnerability and Project Context. Across 60 CVE reproduction attempts, we observed a 63.3% (38/60) overall success rate, with PoC availability in the advisory significantly helping the reproducibility (60% with PoC). Web-centric ecosystems (TypeScript/JavaScript/PHP) showed markedly higher success than compiled/system-heavy stacks (C/C++/Go/C#/Kotlin/Java), where build and environment issues dominated failures. Classic input-validation and injection classes (e.g., SQLi, XSS, path traversal, prototype pollution) reproduced reliably when endpoints were reachable, whereas access-control vulnerabilities underperformed, likely due to the need for realistic authentication/authorization context. The principal blockers were build failures (36%) and timeouts (40%), indicating that improvements in environment provisioning and adaptive budgeting are as critical as exploit logic for scaling automated CVE reproduction.
RQ1: CVE-Genie’s success first depends on whether it can build the vulnerable project, after which performance varies with vulnerability type and language, e.g., web-based and input-validation bugs are easier to reproduce, while compiled/system-heavy stacks and access-control related CVEs are harder.
4.3 Robustness to Loss of CVE Context
Dataset ($D_{R}$). To evaluate CVE-Genie’s robustness to incomplete CVE data, we perform an ablation study using 15 CVEs from $D_{S}$, each with complete information (i.e., CVE description, PoC, patch commit, and security advisory) and reproduced in the first three baseline iterations (Section 4.2). We call this dataset $D_{R}$. We intentionally select these “easier” CVEs because they serve as a controlled probe for ablation: if performance degrades even on CVEs that CVE-Genie can normally handle reliably, it indicates that the missing information is truly critical. As shown earlier, the availability of the PoC is significant. To simulate other incomplete context scenarios we create three variants of $D_{R}$ by systematically removing: (a) the security advisory, (b) the patch commit, and (c) all except the CVE description.
Methodology. For each variant of $D_{R}$, we run CVE-Genie three times per CVE.
Results. Removing the patch commit lowers performance to 10/15 (67%), even when the PoC is available, as it significantly impairs the model’s ability to localize the vulnerability. For example, in CVE-2025-31481, an authorization flaw in the PHP-based API Platform Core, the system lacked direction, explored irrelevant parts of the codebase, and eventually exhausted its cost budget. Removing the security advisory also results in a drop to 11/15 (73%). Most failures here are due to environment setup and verification issues rather than exploit synthesis errors. As security advisories often contain contextual details that guide environment construction and verification strategy. On the other hand, even with only the CVE description, CVE-Genie successfully reproduces 9/15 (60%), particularly for vulnerabilities with recognizable patterns (e.g., SSRF, CSRF). This highlights the model’s baseline robustness under minimal context.
RQ2: CVE-Genie can reproduce CVEs with limited information. Code-based PoCs are most helpful, but in their absence, patch commits offer valuable localization cues. However, reproduction success decreases with minimal context.
4.4 Large Scale Evaluation
Figure 9: Breakdown of CVE-Genie’s run on $D_{L}$
Dataset $(D_{L})$. For our large-scale evaluation, we curate a dataset, $(D_{L})$, comprising all CVEs published between June 2024 and May 2025 that include publicly available source code. Since these CVEs postdate knowledge cutoff of LLMs (o3 and o4-mini) used in CVE-Genie, they ensure zero prior exposure. $D_{L}$ includes 841 CVEs across 186 CWEs, 29 programming languages, and 440 open-source projects. Our statistical analysis shows approximately 55% of CVEs in $D_{L}$ fall under the top 25 most dangerous CWEs [37], and around 49% lack a PoC in their advisory.
Methodology. We use the same methodology as Section 4.2, and run CVE-Genie on $D_{L}$ three times, iteratively.
Results. CVE-Genie reproduced 428 CVEs out of 841, spanning 22 programming languages, 141 vulnerability types, and 267 projects (see Table 4). To better understand outcomes, we manually analyzed 50 successful and 50 failed cases, and categorize the main reasons for successes and failures.
Successes. Web vulnerabilities (XSS, CSRF, SQLi, path traversal) were most reproducible, especially in interpreted languages (Python, JavaScript, PHP, Ruby, TypeScript), as shown in Table 4. Projects with clear setup instructions and code-based PoCs with crashing input had the highest success rates, as runnable PoCs directly guided exploit construction.
Failures. Memory-safety bugs (C/C++), concurrency issues, and UI-dependent flaws were hardest to verify. Many failures were due to brittle builds, poor project documentation, or resource-heavy environments; however, 24% failed from time/cost overruns. PoCs expressed only in natural language contributed little, often leaving too much ambiguity for reliable reproduction. Compiled toolchains (C/C++, Rust, Go) were prone to the most setup problems.
Project # Re CVEs # T25 CWEs # Oth CWEs LoC Lang
lunary-ai 35 13 22 ~70k TypeScript
WeGIA 17 15 2 ~1M PHP
cpython 10 2 8 ~1.8M Py, C
vite 6 6 0 ~107k TypeScript
yeswiki 6 6 0 ~234k PHP
symfony 5 4 1 ~1.6M PHP
llama_index 5 4 1 ~1.1M Py
vim 5 3 2 ~1.2M Vim Script, C
directus 5 2 3 ~398k TypeScript
rails 4 2 2 ~450k Ruby
phpipam 4 2 2 ~225k PHP
assimp 4 4 0 ~562k C++
siyuan 3 1 2 ~265k Go, TypeScript
cvat-ai 3 3 0 ~320k Py, TypeScript
MobSF 3 1 2 ~132k JS, Py
mlflow 3 1 2 ~870k Py
Table 4: Top 15 projects identified in the large-scale evaluation of CVE-Genie, showing the number of reproduced CVEs, their classification into CWE Top 25 vs. other CWEs, lines of code (LoC), and primary languages.
RQ3: At scale, CVE-Genie reproduced 428 / 841 CVEs, spanning 267 projects, 22 programming languages, and 141 CWE types, demonstrating broad applicability across ecosystems (see Table 4).
4.5 Importance of CVE-Genie’s Design
Ideal Case Knw Build Pre- Reqs Feed-back Critic Agents Single Agent
15 / 15 9 / 15 13 / 15 5 / 15 8 / 15 0 / 15
- $\uparrow$ 30% Exploit Failure $\uparrow$ 27% Max Tool $\downarrow$ 67% Reprod Rate $\uparrow$ 47% False Reprod -
Table 5: Ablation study of CVE-Genie’s design.
Dataset. For this ablation study, we use the version of $D_{R}$ dataset which contains all essential information fields for 15 CVEs. As established earlier in Section 4.3, omitting these critical fields can affect accurate CVE reproduction. Therefore, in this ablation, we ensure all CVEs retain their complete critical information. This allows us to isolate the effect of changes to the design of CVE-Genie. If performance degrades when a specific change is made, we can attribute that drop solely to that change, since the CVE inputs remain otherwise fully informative.
Methodology. As Data Processor, Setup Developer, Exploit Developer, and Verifier Developer are the four indispensable components of CVE-Genie for end-to-end CVE reproduction, for this evaluation of CVE-Genie’s design we use the following five ablation settings:
  1. 1.
    No Knowledge Builder: Bypass the Knowledge Builder and feed raw data directly to all agents.
  2. 2.
    No Pre-Requisite Developer: Eliminate the Pre-Requisite Developer, allowing the Setup Developer to independently plan and execute the entire environment setup from scratch.
  3. 3.
    No Feedback Loops: Enforce single-shot execution without iterative refinement, by removing all feedback loops.
  4. 4.
    No Critics: Remove the Verifier Critic and treat a CVE as reproduced if the final scripts pass the Flag Checker.
  5. 5.
    Single Monolithic Agent: Combine all modular agents into a single agent with access to same tools, thereby evaluating the performance of a standalone LLM without CVE-Genie’s agentic design and structured guidance.
Results. In our ablation study (Table 5), we find that removing the Knowledge Builder drops reproduction to 9/15, as agents struggled with unstructured advisories and lost context, causing 30% more exploit failures. Eliminating the Pre-Requisite Developer had a mild impact (13/15), but overloaded the Setup Developer, which lead to 27% more tool-call limits being hit, particularly on complex repositories such as CVE-2025-32389 in Nameless, a website software for Minecraft servers. Feedback loops proved most critical: without them, performance collapsed to 5/15 (–67%). On the other hand, dropping critic agents cut results to 8/15 and raised false reproductions by 47%. Finally, collapsing into a single agent yielded 0/15 reproductions, even advanced standalone LLMs (o3) failed, underscoring the necessity of CVE-Genie’s modular, feedback-driven, critic-guided design.
RQ4: Standalone LLMs are currently unable to reproduce CVEs end-to-end. CVE-Genie’s modular design remains the most effective approach for enabling LLMs to perform this task comprehensively (Table 5).
5 Limitations and Future Work
Vulnerabilities that Involve UI Interactions. Currently, CVE-Genie only supports command-line interface (CLI) interactions. However, many web-based vulnerabilities require interaction with a graphical UI, making them difficult to trigger via CLI. Future work should explore integrating CVE-Genie with UI environments such as web browsers.
Multimodal CVE Knowledge Curation. Due to lack of image and video processing support, CVE-Genie struggles with CVEs where the PoC is provided as an image or video. Future work can address this limitation by integrating CVE-Genie with multimodal LLMs. Additionally, by leveraging external CVE resources through LLM-based deep search tools, beyond standard vulnerability databases, may also help improve reproduction success rates [38].
Over Critique. As shown in Figures 5(b), 6(b), and 7(b), all critics exhibit a lower TPR compared to their TNR. While this helps reduce false positives, it also results in missing some valid reproductions, particularly in cases where critics requested detailed setup, exploit, or verifier information that the developer could not sufficiently provide.
Cost. Running CVE-Genie costs approximately \$2.77 per CVE. While open-source models do not perform well and their large-scale deployment remains a challenge. To support further research, we release logs from all runs as a novel dataset to help the community fine-tune and improve local models’ performance for CVE reproduction.
6 Conclusion
In this paper we presented CVE-Genie, an automated framework designed to reproduce real-world vulnerabilities at scale. This has considerable benefit for downstream security research (see Appendix B), which suffers from a shortage of diverse high-quality data suitable for research and development of new tooling for automated vulnerability discovery and repair. CVE-Genie successfully reproduced around 51% (428) of 841 CVEs, with on average \$2.77 cost per CVE, across a vast variety of projects, CWEs, and programming languages. To the best of the authors’ knowledge, no other framework is capable of such comprehensive results at this scale and efficiency.
Ethical Considerations
The goal of our work is to support defenders and solution developers by accelerating vulnerability triage and tool development for vulnerability detection and patching. CVE-Genie, which does not create or determine new vulnerabilities, provides a powerful framework capable of providing recreations of CVE reports to provide up-to-date and reproducible CVE environments as benchmark test cases. These benchmarks can then be used to rigorously test and enhance vulnerability detection and patching systems, thereby understanding their performance and ideally accelerating their response capabilities. Following the beneficence criterion of the Menlo Report [14], we are confident that our system provides more benefits than potential harms to the security community and to society.
Open Science
We will open source CVE-Genie’s source code, logs for all experiments in Section 4 including all agent conversations, and our dataset of 428 reproduced CVEs.
Appendix A CVE-Genie Specifications
CVE-Genie is built on LangChain and supports easy design and integration of tools for LLMs. For CVE-Genie to effectively reproduce CVEs, it require access to a project’s source code. However, due to LLMs’ limited context windows, large project files cannot be processed all at once. To address this, we developed primitives enabling directory browsing using “project directory tree” and command execution:
  1. 1.
    get_file: The LLM agent specifies the file path and the number of lines it wants to read from a given offset. To prevent context overflow and improve efficiency, it reads a maximum of 300 lines at a time. If more content is needed, the agent can scroll up or down to access it.
  2. 2.
    write_to_file: The LLM provides the path and content for a file, which we then write to the specified file. This facilitates modifications and updates within the project.
  3. 3.
    execute_ls_command: The LLM specifies a directory, and we execute the ls command within it, returning the output. This command is distinct from execute_linux_command due to its major role in project directory exploration and its frequent use.
  4. 4.
    execute_linux_command: The LLM issues Linux commands to be executed from the project’s root directory. Foreground commands have a 300-second timeout, with standard output and error logged to separate files. After completion, we return up to the last 100 lines of the log and the log file path which LLM can read using get_file for further analysis. Background execution is supported for non-blocking tasks (e.g., starting a web server). For these, we show output after 5 seconds, provide the log file path, and indicate whether the process is still running.
  5. 5.
    set_environment_variable: This tool allows for setting environment variables for subsequent execute_linux_command calls. We observed that LLMs frequently issue calls to export environment variables, which are not retained in successive runs as each command is executed in a separate shell. To address this, we enable the LLM to set these variables, which are then passed to commands. There is also an option to clear all set variables.
Restrictions. Each tool-calling agent is limited to 60 tool calls. This limit was found to be optimal in preliminary experiments for maximizing CVE reproduction across all LLMs.
Output Formatting. For all agents requiring output in a specific format, we add a gpt-4o-mini-based format corrector to the output parser. If parsing fails, the raw output is passed to gpt-4o-mini for formatting. If formatting fails after three attempts, an error is returned.
Appendix B Applications of CVE-Genie
Category Application Benefits brought by CVE-Genie
Vul.
detection
ML-based detection Training and testing data
Dynamic fuzzing Benchmarks and initial fuzzing seeds
Static analysis Benchmarks and static rules
Vul.
triage
Root cause analysis Enrich inputs and enable more tools
Exploit generation Ingredients for exploit chaining
Vul.
Patching
Patch generation Benchmarks and testing environments
Patch verification Inputs and environments
Other
applications
LLM secure code generation Benchmarks and testing environment
Penetration testing Construct pen. test tasks and environment
Attack detection Benchmarks and detection rules
Table 6: Tasks that benefit from CVE-Genie.
As shown in Table 6, reproducible CVEs can be used for a plethora of security analysis tasks, including vulnerability detection, triage, exploitation, and patching.
Vulnerability Detection. Most existing vulnerability detection datasets are built using patch commits from CVEs (see Table 1). They typically label the pre-patch version of a function as vulnerable and the post-patch version as benign [40, 15, 7]. However, extracting isolated functions often strips away critical context [51], and functions labeled for a single vulnerability may contain others, resulting in noisy labels. Using reproducible CVEs with working PoCs can address these issues. By executing PoCs during function code extraction, we can ensure that vulnerable execution traces are included, preserving necessary context and confirming the presence of vulnerabilities. Additionally, this approach supports both static and dynamic analysis, unlike most existing datasets (e.g., SVEN [21], BigVul [15], PrimeVul [13]), which are limited to static methods.
Vulnerability Triage and Patching. Reproducible CVEs are essential for effective vulnerability triage. They help eliminate false positives and provide an executable environment with known vulnerable inputs, which aids in accurately analyzing and reproducing bugs. This capability enables advanced techniques like reverse execution and backward taint analysis for root cause identification. CVE-Genie supports this process by offering exploits for individual vulnerabilities, which can be combined into complex exploit chains for studying multi-stage attacks. It also serves as a benchmark for evaluating patching methods, providing rich contextual data such as crash stack traces and outputs. Additionally, its PoCs assist in verifying and refining patches.
LLM Insecure Code Generation. Reproducible CVEs are valuable for evaluating the security of code generated by LLMs. Recent studies have shown that LLMs can produce insecure code, particularly in security-critical contexts [59, 55]. Various benchmarks evaluate secure code generation capability of LLM, such as, CyberSecEval [55] prompts an LLM to generate code snippets and uses rule-based detectors to identify insecure code, however, rule-based detection often leads to false positives. To improve this, SecCodePLT [59] leverages CWE data and manually crafts dynamic test cases to evaluate secure code generation. Reproducible CVEs offer a scalable alternative because these CVEs are confirmed to be vulnerable, they can be used to assess LLMs’ ability to generate secure code. By prompting the LLM to implement the functionalities in the vulnerable functions, and using the associated PoCs to test the output, we can determine whether the generated code is vulnerable.
Penetration Testing and Attack Detection. CVE-Genie offers a valuable resource for both penetration testing and cybersecurity training. It provides ready-to-use exploitation scenarios along with verified solutions, making it ideal for educating future penetration testers and developing automated testing tools. For defenders, these exploits can be analyzed to identify attack patterns, which can enhance detection mechanisms such as intrusion detection systems (IDS) or malware classifiers.