### v2

# From CVE Entries to Verifiable Exploits: An Automated Multi-Agent Framework for Reproducing CVEs

Abstract
High-quality datasets of real-world vulnerabilities and their corresponding verifiable exploits are crucial resources in software security research. Yet such resources remain scarce, as their creation demands intensive manual effort and deep security expertise. In this paper, we present CVE-Genie, an automated, large language model (LLM)-based multi-agent framework designed to reproduce real-world vulnerabilities, provided in Common Vulnerabilities and Exposures (CVE) format, to enable creation of high-quality vulnerability datasets. Given a CVE entry as input, CVE-Genie gathers the relevant resources of the CVE, automatically reconstructs the vulnerable environment, and (re)produces a verifiable exploit. Our systematic evaluation highlights the efficiency and robustness of CVE-Genie’s design and successfully reproduces approximately 51% (428 of 841) CVEs published in 2024-2025, complete with their verifiable exploits, at an average cost of \$2.77 per CVE. Our pipeline offers a robust method to generate reproducible CVE benchmarks, valuable for diverse applications such as fuzzer evaluation, vulnerability patching, and assessing AI’s security capabilities.
1 Introduction
Tens of thousands of software vulnerabilities are discovered every year, with more than 40,000 vulnerabilities tracked by the National Vulnerability Database (NVD) [42] in 2025 alone. Many of these vulnerabilities are cataloged using the CVE format, which provides a standardized identification system to help organizations address security flaws in both software and hardware. Still, existing vulnerability detection tools, such as those listed by OWASP [44], struggle to keep up with the increasing volume of code released each year. As a result, many vulnerabilities remain undiscovered in codebases for years (e.g., up to 2 years in Chromium and 7 years in OpenSSL [3]) posing significant security risks.
Developing effective automated vulnerability detection techniques depends on high-quality datasets with reliable ground truth, as low-quality data can lead to unreliable evaluations [51] and to learn spurious correlations, especially in machine learning (ML) based approaches [4, 49]. Although public repositories, such as the NVD, include hundreds of thousands of CVEs, they often lack critical details, including vulnerable code, environment setup, working exploits, and validation steps, which are all required to effectively reproduce a vulnerability and to build ground truth datasets for research [38]. This is because the main purpose of vulnerability advisories is to alert system administrators about software that needs to be updated, and concrete exploit code is often kept confidential for ethical reasons. Unfortunately, generating this missing information from CVEs alone is labor-intensive and requires significant expertise, e.g., Mu et al. [38] spent over 3,600 hours with 43 security expertise to reproduce only 368 memory vulnerabilities in Linux.
Figure 1: CVE-GenieOverview.
To address these challenges, the security community has adopted several approaches: manual annotation [21, 65], inserting bugs into real-world code [41, 36], and heuristic-based mining [13, 10, 40, 15, 7]. While these methods have advanced the field, they come with limitations, such as limited support for software and Common Weakness Enumerations (CWE) categories, mislabeled vulnerabilities [50], and data becoming stale. For instance, the DARPA Cyber Grand Challenge dataset from 2016 [11] quickly became outdated due to its small size and low-complexity bugs [20]. Meanwhile, automated data collection helps address some of these limitations, but often tends to be limited in scope, e.g., ARVO [34] automatically curates reproducible builds with triggering inputs from OSS-Fuzz reports, but its scope is limited to memory corruption bugs in C/C++. This highlights a clear need for a dataset curation method that enables fully reproducible vulnerability instances across a wide range of software and vulnerability types.
Recently, LLMs have demonstrated remarkable capabilities in solving complex software engineering (SWE) and security tasks, ranging from automated bug detection [39, 16] and repair [46, 57, 24], to writing fuzzer test harnesses [12]. Moreover, the shift to agentic workflows [63, 8, 62, 32], in which LLM-powered agents are provided access to real-world tools [1], has further pushed the capabilities of general-purpose LLMs and led to significant performance gains on extremely difficult real-world SWE-related benchmarks, such as SWE-bench [25]. This raises the question of whether LLM agents could also be useful in automatically setting up and reproducing existing CVEs, allowing us to build reliable, large, and realistic benchmarks for the security community.
In this paper, we present CVE-Genie, a fully-automated, LLM-driven, multi-agent framework for end-to-end CVE reproduction. CVE-Genie instantiates what we define as key properties of an ideal CVE reproduction system, captured by the acronym EAGER. Specifically, an ideal system: Generates working Exploit/proof-of-concept (PoC); Builds Assessors/verifiers for the exploit; Generalizes across CWEs, languages, and project types; Is End-to-end automated; and Rebuilds vulnerable environments. Specifically, this paper makes the following contributions:
  1. 1.
    We develop CVE-Genie, an automated end-to-end CVE reproduction pipeline that extracts CVE data (e.g., software vulnerable version source code, advisories, patches, cwes, etc.), rebuilds the vulnerable environment, and generates exploits and verifiers to assess the reproduction.
  2. 2.
    We design CVE-Genie to demonstrate compositional intelligence by a novel architecture, systematically selected LLMs and prompts, and reliable EAGER-style CVE reproduction. As a result, CVE-Genie achieves strong performance across diverse CVEs, vulnerability types, programming languages, and projects, even when CVE information is incomplete.
  3. 3.
    We systematically evaluate CVE-Genie’s architecture and demonstrate that each component is essential for optimal performance. Notably, even advanced LLMs like o3, when used standalone, are incapable of performing EAGER-style reproduction for even a single CVE.
  4. 4.
    We ran CVE-Genie on 841 CVEs (published between June 2024 and May 2025) and successfully reproduced 428 CVEs across 267 projects, 141 CWEs, and 22 programming languages, highlighting our generalizability.
  5. 5.
    We make our framework, source code, datasets of reproduced CVEs, and full logs of agent interactions publicly available, providing an ongoing valuable contribution for downstream research tasks (see Appendix C and  F).
2 Background and Related Work
Name Source
Real
# Projects
# CWEs
# Lang
Automated
Reproducible
PoC
Verifier
SARD [41] Manual
~120 ~150 6
SVEN [21] Manual
n/a 9 3
Devign [65] Manual
4 n/a 2
SecLLMHolmes [54] Manual
4 8 3
VulChecker [36] Manual
n/a 5 2
Draper [52] SA tool
n/a 9 2
D2A [64] SA tool
6 7 2
VUDENC [56] Sec. issue
812 7 1
DiverseVul [10] Sec. issue
797 150 12
PrimeVul [13] Patch diff
755 140 2
BigVul [15] Patch diff
348 91 2
CrossVul [40] Patch diff
1,675 168 40
CVEfixes [7] Patch diff
1,754 180 30
ARVO [34] OSS-Fuzz
273 4 2
CVE-Bench [66] Public CVEs
36 8 6
Mu et al. [38] Public CVEs
1 4 1
CVE-Genie Public CVEs
267 141 22
Table 1: Overview of vulnerability benchmark datasets. Real: synthetic , real-world , mixed . CVE-Genie results reflect CVEs from Jun 2024–May 2025.
2.1 Existing Benchmarks and Limitations
To evaluate various vulnerability detection techniques (rule-based methods [44, 58], ML-approaches [36, 30, 31, 29], and LLM-based systems [17, 19, 47, 18]), the community needs standardized benchmarks. Prior work has explored four main approaches for addressing these challenges. (1) Manual efforts involve analyzing [21] or reproducing [66] vulnerabilities in real-world code, or crafting synthetic examples [54], but datasets are small, lack ecological validity, and often suffer labeling errors due to human mistakes (e.g., 50% data in Devign [65] have wrong labels  [50]). (2) Automated labeling using static analysis (SA) tools, such as bandit [6] or infer [23], reduces manual effort but introduces many false positives [48, 64]. (3) Mining real-world data, such as GitHub issues or CVE patch commits, assumes that all removed or modified code is vulnerable; However, this assumption often leads to mislabeled samples, as Risse et al. [50] highlighted for DiverseVul [10] and BigVul [15]. (4) Relying on proven sources, such as OSS-Fuzz, offers verified exploits and sanitizer feedback; However, these are restricted to certain bug types, primarily memory corruption in C/C++, and do not generalize across languages or vulnerability classes [34].
Besides data quality, existing datasets also lack necessary artifacts for dynamic evaluation beyond static labels [4, 49, 54], including dynamic execution environments and reproducible exploits, which is difficult to obtain [38]. Table 1 provides an overview of the current state of vulnerability benchmarks.
2.2 Call for EAGER Style for CVE Reproduction
A CVE ID is a unique identifier for publicly disclosed security vulnerabilities in real-world software. Each CVE ID corresponds to a CVE entry that includes key details for that CVE, such as vulnerability description, source code, affected versions, security advisories, CWE classifications, and patches, enabling security professionals to assess and mitigate risks. Creating a dataset from the standard CVE database is a promising way towards addressing the aforementioned challenges (see Appendix F). Therefore, we introduce EAGER, a set of crucial criteria for ideal CVE reproduction:
  • Exploit Generation – (re)creation of an exploit or PoC that reliably triggers the vulnerability in the CVE.
  • Assessment – inclusion of a “verifier” or “sanitizer” capable of assessing whether the generated exploit or PoC successfully triggers the vulnerability.
  • Generalization – reproduction of CVEs across diverse CWEs, programming languages, and software projects.
  • End-to-end Automation – execution of all stages of CVE reproduction in a fully automated manner.
  • Rebuild project – reconstruction of the original vulnerable environment to facilitate exploit/PoC execution.
However, none of the existing methods satisfy all the criteria (see Table 2). For example, manual effort for producing EAGER datasets does not scale and is not generalizable [38, 66]. ARVO [34] focuses solely on memory-related bugs from OSS-Fuzz C/C++ projects, limiting its scope. In contrast, CVE-Genie is the only framework that satisfies all criteria of EAGER reproduction.
Method Exploit Assess General E2E Rebuild
Manual [32, 66]
Fuzzers
Metasploit [35]
ARVO [34]
CVE-Genie
Table 2: Comparing Potential Vulnerability Reproduction Methods and their EAGER attributes.
2.3 Large Language Models and Agents
LLMs are transformer-based neural network models with billions of parameters. They have demonstrated remarkable performance across various application domains, including question answering [43], mathematical reasoning [2], and code generation [9, 5]. Recent work also shows their utility in security tasks, e.g., bug detection and repair [46, 57, 24] and writing fuzzer test harnesses [12].
LLM agents are systems that combine LLMs with tools. They can finish complex tasks via a sequence of actions, including planning, generation, and tool executions [53, 22, 45, 27, 26]. An agentic system can have multiple agents, each with a sub-task, such as reasoning, critiquing, generating, perceiving, remembering, or teaching, as well as its own workflow and tool sets. In security, researchers have started to construct LLM agents to find and repair vulnerabilities. There are multi-agent patching systems, where different sub-agents are responsible for fault localization, patch generation, and validation (e.g., PatchPilot [28], and PatchAgent [61]). Through the recent DARPA AIxCC competition, researchers constructed agentic-based systems with end-to-end capabilities for vulnerability detection and patching [18]. These recent successes provide the argument for using LLM agents in vulnerability analysis and software development, as well as for building an EAGER-style automated end-to-end CVE reproduction framework.
3 CVE-Genie
We design CVE-Genie to exhibit compositional intelligence, i.e., the ability to solve complex tasks by decomposing them into meaningful sub-tasks whose outputs are explicitly composed and verified to form a correct end-to-end solution. To this end, first, we design a novel architecture that decomposes CVE understanding and reproduction into four interdependent modules: (1) the Processor (Section 3.1), which retrieves source code and constructs a structured knowledge base; (2) the Builder (Section 3.2), which reconstructs the vulnerable environment; (3) the Exploiter (Section 3.3), which reproduces the exploit; and (4) the CTF Verifier (Section 3.4), which generates a verifier for the exploit. Second, rather than relying on a single LLM or a fixed prompting strategy across tasks, we systematically select the most suitable LLMs and engineer task-specific prompts for each component of CVE-Genie, resulting in an optimized end-to-end configuration (Section 3.5). Third, we bridge the gap identified in Section 2, namely the lack of any system capable of true EAGER-style CVE reproduction, and ground CVE-Genie’s design in the following guiding principles:
  1. 1.
    Modular Task Decomposition: LLMs struggle with long, complex contexts [54] and tasks [25]. Thus, CVE-Genie decomposes reproduction into focused modules and sub-modules handled by specialized agents with systematically engineered prompts, providing the most suitable architecture for end-to-end CVE reproduction (Section 4.2).
  2. 2.
    Robustness to Incomplete Data: Since missing CVE fields can lower reproduction success by up to 44% [38], CVE-Genie mitigates this by relying on patch or source-code analysis when advisories/PoCs are unavailable (Section 4.3).
  3. 3.
    Reliability through Self-Critique: To counter LLMs’ reasoning limits [54], each module uses paired developer and critic agents in a ReAct-style loop [60], enabling iterative refinement and internal verification (Section 4.2).
Together, these principles elevate CVE-Genie beyond a fixed engineering pipeline and enable CVE-Genie to operationalize compositional intelligence in a fully automated setting, yielding the first practical, generalizable, and end-to-end framework for EAGER-style CVE reproduction (see Appendix E for further details). The following subsections detail the workings of each CVE-Genie component, using CVE-2024-4340 as a running example corresponding to the end-to-end reproduction shown in Figure 2. Moreover, we provide a detailed case study on CVE-2024-5129, a logic vulnerability of missing authorization in lunary-ai, in Appendix H.
Figure 2: CVE-Genie architecture and an end-to-end example of workflow of reproduction for CVE-2024-4340, i.e., Denial of Service due to RecursionError in sqlparse < v0.5.0. See artifact of CVE-2024-4340 complete reproduction run here - https://github.com/BUseclab/cve-genie/tree/main/results/CVE-2024-4340
3.1 Processor
This module comprises two sub-modules: the Data Processor, which locates the vulnerable project’s source code and gathers relevant raw resources, and the LLM-based Knowledge Builder, which transforms these resources into a structured knowledge base. Below, we provide detailed functionalities of these sub-modules.
3.1.1 Data Processor 1
For the given CVE ID, this sub-module identifies and extracts the vulnerable version of the software along with all relevant CVE details, as follows:
Source Code Extraction. The Data Processor begins by locating the source code for the vulnerable project linked to the CVE. It searches the GitHub URLs referenced in the “cvelist” 111cvelist” is an automated pilot program for CVE submission through GitHub (https://github.com/CVEProject/cvelist) to identify the project repository. For closed-source software, Data Processor allows users to supply the repository URL. Once the source code is located, it identifies affected software configurations: vulnerable versions, platforms, or settings. For instance, CVE-2024-4340 affects all versions of “sqlparse” prior to v0.5.0. Accordingly, the latest affected version (v0.4.4) is retrieved using the GitHub API, and its source code is downloaded (as shown in Figure 2).
Vulnerability Information Extraction. After the vulnerable version of the project source code is downloaded, Data Processor extracts the following four key pieces of information from “cvelist”, if available: (1) CVE description: A high-level summary of the vulnerability in the target project. (2) CWE data: CWE information, a categorization system for hardware and software weaknesses that associates the CVE with a specific vulnerability type, e.g., CWE-674 (Uncontrolled Recursion). (3) Patch commits: The git-diff of code changes in the target source code that fix the vulnerability. These commits help LLMs localize the issue, understand its root cause [21, 13, 15], and generate effective exploits. (4) Security advisories and PoC: We filter URLs referenced in the CVE entry by keywords (e.g., “security”, “advisory”, “bounty”, etc.), and scrape their contents.
Output. The vulnerable version of a project’s source code alongside its directory structure and CVE raw data.
3.1.2 Knowledge Builder 2
This agent is instructed to analyze, distill, and organize the raw information, extracted by the Data Processor, into a structured knowledge base while retaining essential details required to reproduce the given CVE. This includes extracting PoC or exploit instructions from security advisories, which are invaluable for accurately reproducing the vulnerability. For instance, CVE-2024-4340 has a PoC provided in its security advisory, and the patch commit helps understand the root cause by highlighting where/how the vulnerability was mitigated. As shown in Figure 2, the Knowledge Builder included both of these details in the knowledge base. In case an advisory does not include an exploit, the LLM agent is tasked with generating an overview of what an exploit might entail. Additionally, capturing the patch details helps to localize the root cause of the vulnerability [21, 13, 7] and aids in crafting an effective exploit. This knowledge base acts as the long-term memory for agents, helping them to effectively reproduce CVEs, by including only essential information in their context to avoid overloading the LLM’s context window.
Output. A structured CVE knowledge base providing details of CVE, including its associated CWEs, affected software configurations, root causes (if provided in the security advisory), and essential information from security advisories and patch commits.
3.2 Builder
Once the knowledge base is populated and the vulnerable version of the project’s source code is downloaded, the next objective is to build the project in a way that allows the exploit to be executed. To achieve this, our pipeline employs two developer LLM agents: the Pre-Requisite Developer Agent, which analyzes and plans the environment setup, and the Setup Developer Agent, which executes this plan to configure the vulnerable environment. As well as one critic, the Setup Critic Agent, which analyzes the logs of the Setup Developer and evaluates the setup for the project, allowing the system to correct its mistakes. Below, we discuss these in detail.
3.2.1 Prerequisite Developer Agent 3
We introduce this agent as an initial exploration and planning step before actually setting up the vulnerable project. This is done for three key reasons.
(1) Our preliminary experiments reveal that many projects contain inaccurate and incomplete information in their README files (e.g., vulnerable version v1.2.7 of lunary for CVE-2024-5129, includes incorrect paths to .env files that are crucial for the setup). Without the Prerequisite agent, these details were consistently overlooked by LLMs, causing setup failures. Identifying and correcting such issues early reduces unsuccessful attempts, conserves the LLM’s context window, and improves the likelihood of successful setup.
(2) During environment setup, the agent’s job is to explore and analyze the source code to identify critical components and to build the project by executing a series of commands. For large projects, having to do both tasks with one agent can exceed the context limit and reduce efficiency (as demonstrated in the ablation study in Section 4.2). To address this, we delegate the exploration of the code base to a dedicated agent, the Pre-Requisite Developer, allowing the setup process to focus more on building the project.
(3) The Pre-Requisite Developer also defines the “expected state,” i.e., when the project is fully set up and ready for exploitation. This state guides the Setup Developer Agent in verifying the correct configuration of the vulnerable environment. For example, in CVE-2024-4340 (Figure 2), the Pre-Requisite Developer explored the root directory and key files such as README and sql.py to understand the project context, then specified the expected state as having sqlparse version v0.4.4 installed for the exploit to work correctly.
Tools. For this step, we only provide the LLM agent with access to read-only tools, i.e., execute_ls_command, and get_file_by_name, which the agent can use to traverse the code base and read files, because we do not want the agent to execute any commands or write to files in this phase (see Appendix E for tools details).
Output. (1) Detailed overview of the project, (2) important files to pay attention to during the setup, (3) required services and their configurations, and (4) expected state of the project upon successful setup to facilitate verification.
3.2.2 Setup Developer Agent 4
The Setup Developer begins in the directory containing the vulnerable version’s source code and focuses on configuring the project based on instructions from the Pre-Requisite Developer, exploring the codebase when necessary. For CVEs involving third-party libraries, our initial experiments showed that setup is greatly simplified by using package managers like pip or npm to install specific vulnerable versions instead of building from source. For example, in the case of CVE-2024-4340 affecting sqlparse, the Pre-Requisite Developer instructed the installation of version 0.4.4, and the Setup Developer executed pip install sqlparse==0.4.4 directly (command 4a, Figure 2), resulting in a smooth setup. If the package manager fails, the Setup Developer falls back to building from source. Typically, the Pre-Requisite Developer also recommends running a basic PoC to confirm readiness before handing off to the next agent (e.g., commands 4e and 4f in Figure 2).
Tools. For this step, we allow access to the following tools, i.e., execute_ls_command, execute_linux_command, get_file, write_to_file, and set_environment_variable.
Output. This agent returns a final decision whether the setup is successful or not. If successful, the output also includes instructions on how another agent can access the running project.
3.2.3 Setup Critic Agent 5
The Setup Critic evaluates whether the project setup performed by the Setup Developer is correct and complete, ensuring that the environment is properly configured for the vulnerability described in the CVE knowledge base to be exploited. Additionally, this agent plays a critical role in detecting deceptive or shortcut behaviors by the LLM, such as fabricating a mock project instead of performing a genuine setup, or deliberately injecting vulnerabilities into the codebase to simplify exploitation. These undesirable patterns are discussed in detail in Section 3.5.1, which informs the instruction design for the Setup Critic. Based on this, the critic is equipped to identify both functional and security-related flaws in the setup process.
Output. A comprehensive analysis of the setup logs, a binary decision on whether the setup is valid and complete, and actionable feedback for correcting issues or improving the setup if necessary.
1"""
2PoC - CVE-2024-4340 (sqlparse < 0.5.0)
3EXAMPLE THAT CRASHES v0.4.4
4    python3 exploit.py 10000
5"""
6import sys
7import sqlparse
8def main() -> None:
9 if len(sys.argv) != 2:
10 print("Usage: python3 exploit.py <depth>")
11 sys.exit(1)
12 arg = sys.argv[1]
13 try:
14 depth = int(arg)
15        payload = "[" * depth + "]" * depth
16 except ValueError:
17 payload = arg
18    sqlparse.parse(payload)
19if __name__ == "__main__":
20 main()
(a) Exploit script for CVE-2024-4340. It constructs a deeply nested list payload and submits it to sqlparse.parse(), triggering uncontrolled recursion. The included comment documents a crashing input (depth 10,000) that reliably triggers the vulnerability.
1Traceback (most recent call last):
2 File "poc.py", line 20, in <module>
3 sqlparse.parse(payload)
4 ....
5 "sqlparse/sql.py", line 214, in flatten
6 yield from token.flatten()
7 [Previous~line~repeated~983~more~times]
8RecursionError: maximum recursion depth exceeded
(b) Execution traceback produced by the exploit, showing a RecursionError after excessive recursive calls within sqlparse.
Figure 3: PoC for CVE-2024-4340 in sqlparse v0.4.4, showing the exploit script and the resulting crash.
3.3 Exploiter
After successfully configuring the vulnerable system, the next step in our pipeline is to generate an exploit for the given vulnerability. CVE-Genie accomplishes this using two LLM agents: the Exploit Developer, which develops the PoC for an exploit in the vulnerable environment, and the Exploit Critic, which analyzes the logs of the Exploit Developer and evaluates the exploit for the given vulnerability. Below, we discuss their detailed functionalities.
3.3.1 Exploit Developer Agent 6
The Exploit Developer is responsible for crafting and demonstrating exploits for vulnerabilities listed in the CVE knowledge base on pre-configured vulnerable systems. If a PoC is provided, the agent replicates and verifies it by executing the script with the appropriate triggering input before delivering the final, working PoC. For example, in CVE-2024-4340 (Figure 2), the PoC was available, so the agent validated the setup (6b), demonstrated the exploit (6c) by triggering RecursionError in sqlparse/sql.py (see Figure 3(b)), and submitted the verified PoC script (see Figure 3(a)). If no PoC or exploitation steps are available, the Exploit Developer iteratively analyzes the codebase using available tools to develop a functional exploit. Upon success, it produces a Python PoC script that accepts the crashing input via the command line, includes comments on the expected input format, and provides an example input. This ensures that the CTF Verifier can reliably validate the exploit without repeating the analysis.
Tools. This agent uses the same tools as the Setup Developer.
Output. The agent returns its final decision whether it considers the exploit is successful or not. If successful, it provides an overview of the exploit and a Python PoC script.
3.3.2 Exploit Critic Agent 7
Similar to the role of the Setup Critic, the Exploit Critic is tasked with evaluating the behavior of the Exploit Developer by analyzing its execution logs. Its primary goal is to determine whether the generated exploit is both valid and high-quality. This includes verifying that the exploit could plausibly succeed in a real-world setting; that it does not rely on manipulating or altering the original setup in illegitimate ways; and that it avoids fabricated shortcuts, oversimplified logic, or artificial assumptions that would not hold in practice.
To perform this assessment, the Exploit Critic leverages a predefined set of behavioral patterns and common failure cases, as detailed in Section 3.5.2. By comparing the exploit’s behavior against these known issues, the critic can detect actions that are unfaithful to the original intent or technically incorrect. When flaws are identified, the Exploit Critic provides structured, actionable feedback to improve the robustness of the exploit in subsequent iterations.
3.4 CTF Verifier
After generating a working exploit for a given CVE, the final step is to create a verifier script that confirms whether the exploit was successful. This approach is inspired by security Capture-The-Flag (CTF) challenges, where exploiting a vulnerability reveals a hidden “flag” i.e., a string of random characters or numbers. The CTF Verifier uses three sub-modules to develop and validate such verifiers. The Verifier Developer agent module generates a candidate verifier, the Flag Checker module ensures that the exploit correctly retrieves a flag, and the Verifier Critic agent assesses the quality and effectiveness of the verifier for the given exploit.
1import os, subprocess, sys
2FLAG = "3xploit66full"
3
4def pre_setup():
5    # Configuring PYTHONPATH
6
7def run_exploit():
8 cmd = [sys.executable, "exploit.py", "10000"]
9    return subprocess.run(cmd, timeout=20)
10
11def post_exploit(proc):
12 output = (proc.stdout or "") + (proc.stderr or "")
13    return proc.returncode != 0 and
14            "RecursionError" in output
15
16if __name__ == "__main__":
17 pre_setup()
18 try: result = run_exploit()
19    except subprocess.TimeoutExpired: sys.exit(1)
20 if post_exploit(result):
21 print(FLAG)
(a) Attempt 1 – Weak verifier script: pre_setup() lacks a check for the vulnerable sqlparse version. The exploit is run with a 20s timeout, risking missed RecursionErrors if they occur after the limit. The verification relies solely on detecting the string RecursionError in output, which can be trivially spoofed.
1import os, sys, importlib, traceback
2FLAG = "3xploit66full"
3
4def pre_setup():
5    assert sqlparse.__version__ < "0.5.0"
6
7def run_exploit_inprocess():
8 sys.argv = ["exploit.py", "10000"]
9 try:
10        importlib.import_module("exploit").main()
11 return "no_exception", ""
12    except RecursionError:
13        return "recursion_error", traceback.format_exc()
14 except:
15 return "other_exception", traceback.format_exc()
16
17def exploit_succeeded(status, tb):
18    return status == "recursion_error" and
19            "sqlparse/sql.py" in tb
20
21if __name__ == "__main__":
22 pre_setup()
23 status, tb = run_exploit_inprocess()
24 if exploit_succeeded(status, tb):
25 print(FLAG)
(b) Attempt 2 with Verifier Critic’s feedback – Improved verifier: Validates project version sqlparse < 0.5.0, runs the exploit in-process to avoid timeout issues, and explicitly checks for a genuine RecursionError with traceback originating from sqlparse/sql.py, preventing spoofed or misleading outputs.
Figure 4: Verifier scripts for exploit (in Figure 3) for CVE-2024-4340 corresponding to the run in Figure 2, illustrating the progression from a weak to a robust verifier based on Verifier Critic feedback.
3.4.1 Verifier Developer Agent 8
Following the CTF methodology, we prompt the Verifier Developer to create a general CTF-style Python verifier script for a given PoC script. So, if the verifier script runs the PoC and it successfully triggers the vulnerability, it returns this flag (3xploit66ful). We prompt the Verifier Developer to format the verifier script in three structured steps:
  1. 1.
    Pre-Setup: Prepares the necessary inputs for the exploit, sets up the environment, and prepares a flag to return if the exploit correctly triggers the vulnerability.
  2. 2.
    Exploit Execution: Runs the provided PoC script to trigger the vulnerability. We make sure that the agent cannot modify the PoC script to avoid contamination.
  3. 3.
    Post-Setup: Verifies the success of the exploit, and if successful, the script returns the flag.
This verifier architecture provides consistent and reliable exploit validation across a broad spectrum of vulnerability classes. It supports vulnerabilities with directly observable failure signals, such as a RecursionError in CVE-2024-4340 or an AddressSanitizer: heap-buffer-overflow in CVE-2024-10525, as well as higher-level logic flaws. For example, in the case of the missing-authorization flaw in CVE-2024-5129 (Appendix H), the verifier first inserts a legitimate dataset with a known UUID, then executes the exploit by issuing an unauthenticated deletion request targeting that UUID, and finally confirms successful exploitation by verifying that the dataset has been removed from the database.
Tools. This agent follows the same read-only tools as the Pre-Requisite Developer agent, as in this step we don’t want the model to execute anything, and we do all validations using the Flag Checker and the Verifier Critic.
Output. A Python script that acts as a custom verifier for the exploit/PoC of the given CVE.
3.4.2 Flag Checker 9
Once the verifier script is generated for the given exploit, before passing it to the critic agent, the Flag Checker executes the verifier script to ensure that it runs with the exploit script without any errors and produces the expected output, i.e., the correct flag. For instance, figure 2 shows that the correct flag 3xploit66ful is released when executing the verifier scripts generated for CVE-2024-4340. If a script fails due to a runtime error or if the flag is missing from the output, the system assumes that the verifier is incorrect. Feedback is then sent to the Verifier Developer to revise the script. This process is repeated until a maximum of five attempts is reached, with each attempt ensuring functional correctness of the verifier.
3.4.3 Verifier Critic Agent 10
Once the verifier and exploit scripts pass functional validation via the Flag Checker, they are further examined by the Verifier Critic, which performs a similar evaluative role as the other critic agents, focusing specifically on the response produced by the Verifier Developer. It examines whether the verification process was properly carried out, without assumptions, omissions, or unreliable heuristics. Using the behavioral issues identified in Section 3.5.3, the agent flags incomplete or misleading verification strategies and provides feedback on how the verification could be made more robust or accurate. For instance, Figure 4 shows a detailed example of verifiers generated for CVE-2024-4340 with weak and bypassable verification criteria, which were corrected using the Verifier Critic’s feedback.
3.4.4 Store Reproduced CVE 11
If the given CVE passes the Verifier Critic check, CVE-Genie marks it as reproduced and stores the VM snapshot for the vulnerable environment as well as the exploit and verifier scripts. Moreover, CVE-Genie also stores the metadata, such as LLM cost, time spent, and all agents’ conversations.
3.5 Selecting LLMs and Engineering Prompts for CVE-Genie’s Optimal Performance
Upon finalizing the design of CVE-Genie’s architecture, we perform a systematic LLM model and prompt selection process to identify the best performing LLMs and prompt integrations. We evaluate ten state-of-the-art LLMs (Table 3) across CVE-Genie’s three core modules, Builder, Exploiter, and CTF Verifier, using a diverse set of 60 post–knowledge-cutoff CVEs, while simultaneously engineering and refining task-specific prompts for each module.
Model # Params Max Inp Max Out Reas-oning Open Src Cutoff
o3 n/a 200k 100k 05/2024
o4-mini n/a 200k 100k 05/2024
Claude 3.7 Sonnet n/a 200k 64k 11/2024
Claude 3.5 Sonnet 175B 200k 8k 04/2024
Gemini 2.5 Pro n/a 1M 65k 01/2025
Gemini 2.5 Flash n/a 1M 65k 01/2025
Llama 4 Maverick 400B 1M n/a 08/2024
Qwen 3 235B 128k 64k 06/2024
Deepseek V3 671B 64k 8k 07/2024
DeepSeek R1 671B 128k 8k 07/2024
Table 3: Studied LLMs.
Dataset $(D_{S})$. To evaluate the capabilities of LLMs in reproducing vulnerabilities using CVE-Genie, we construct a small-scale, diverse dataset $(D_{S})$ consisting of $60$ CVEs. These CVEs were published after the latest knowledge cutoff date of the LLMs under study (January 2025), as listed in Table 3. To ensure that $D_{S}$ is both representative and diverse, we select 40 CVEs from the top 25 most dangerous CWE [37] categories, and the remaining 20 CVEs are randomly sampled from the other CWE types. The resulting dataset, $D_{S}$, comprises 60 CVEs drawn from 44 different CWE categories, spanning 18 programming languages, and involving 56 distinct software projects.
Methodology. CVE-Genie is composed of four modules. The Processor’s LLM-based component, the Knowledge Builder, only performs a trivial task of summarization, so we assign a lightweight LLM (o4-mini) to it. In contrast, the remaining modules, i.e., Builder, Exploiter, and CTF Verifier, require strong security vulnerability reasoning and SWE capabilities. Thus, for each of these modules, we systematically select the best performing LLM for “developer” and “critic” tasks using the following 5-step procedure:
(1) Candidate Selection: For each module, we identify subsets of LLMs from Table 3 as potential developers and critics, based on prior evidence of their capabilities, as well as time and cost constraints.
(2) Developer Evaluation: Each candidate developer LLM is run for the given module, and its outputs are manually scored by the authors. The highest-scoring LLM is chosen as the developer.
(3) Critic Prompt Design: We analyze common mistakes made by developers and use these insights to design a module-specific critic prompt, enabling the critic to effectively detect those mistakes.
(4) Critic Evaluation: Each candidate critic LLM is evaluated on true positive rate (TPR) and true negative rate (TNR). When TPRs are equal, we prefer the model with the higher TNR, and vice versa. Then then critic with the best trade-off is selected.
(5) Feedback Loop Assessment: We rerun the chosen best developer–critic pair for the given module, and evaluate the correctness of the critic’s feedback and the developer’s capability to correctly address the critic’s feedback.
3.5.1 Builder
In this section, we systematically select the most optimal configuration of LLMs and prompts for the Builder.
(a) Developer Agent
(b) Critic Agent
Figure 5: Builder Optimal LLM Evaluation
Candidate Selection. Setting up a project from scratch is the most time-consuming and complex part of CVE-Genie, as security advisories or CVEs do not provide instructions for setting up the vulnerable project. To address this, we use two developer agents: one for exploring the project and another for setting it up (see Section 3.2). This process requires extensive tool calls and command executions, making it both slow and costly with reasoning-heavy models like o3 or claude-3.7-sonnet, which can cost \$10–20 per CVE. Therefore, we select fast, cost-effective LLMs with sufficient context capacity, o4-mini, claude-3.5-sonnet, gemini-2.5-flash, llama-4-maverick, deepseek-v3, and qwen3, as developer agent candidates. For critic agents, we opt for stronger reasoning models such as o3, o4-mini, claude-3.7-sonnet, gemini-2.5-pro, and deepseek-r1, as the critic’s task only requires one call, and therefore, they incur minimal overhead.
Developer Evaluation. We ran CVE-Genie up to the Setup Developer stage 4 on CVEs from $D_{S}$ and manually evaluated each developer LLM using the following procedure. (1) For each CVE, we first identified the official installation documentation for the software’s vulnerable version and, based on it, established both the expected final build state and the sequence of commands required for installation. (2) We then reviewed the logs of runs where the LLM reported successful setup, comparing the executed commands against the documented installation process to assess alignment and completeness. (3) Finally, we validated the resulting environment: for libraries or binaries, we confirmed the vulnerable version by invoking the corresponding --version (or equivalent) command (e.g., for CVE-2025-1215 we verified that vim v9.1.1096 was installed), while for server-based software we performed a health check to ensure the service was running and accessible at the expected port.
As shown in Figure 5(a), o4-mini emerged as the best developer LLM for the Builder, successfully setting up 32 out of 60 CVEs projects, while making an average of 20 tool calls per CVE. In contrast, claude-3.5-sonnet mostly gave up on the given setup task, and models like gemini-2.5-flash and open-source LLMs were unreliable: Qwen 3 often described setup steps without executing them using tool calls, while gemini-2.5-flash issued tool calls with frequent syntax errors, eventually leading to the failed setup of the given project.
Critic Prompt Design. From analyzing developer LLM setups, we observed three recurring mistakes: (1) when setup fails, the LLM builds a simplified mock-up substitute project instead; (2) for pip/npm packages, it installs the latest version instead of the specified vulnerable one; and (3) when a server is required, it assumes commands like npm run dev succeed without verifying server health. We designed the critic prompt to detect these errors while reviewing Setup Developer logs.
Critic Evaluation. As shown in Figure 5(b), the o3 model achieves the highest TPR and TNR when evaluating Setup Developer’s execution logs. Most of its false negatives involved setups requiring external hardware, which Exploit Developer cannot exploit anyway. Some valid setups were also mistakenly rejected due to limited post-setup checks. Given its strong balance of recall and precision, we choose o3 as the critic model for the Builder.
Feedback Assessment. We evaluated CVE-Genie up to Setup Developer stage 4, using the best-performing developer (o4-mini) and critic (o3) models. Minor issues flagged by the critic (e.g., limited verification) were typically resolved by the developer in one iteration. However, fundamental issues (e.g., mock-up version of the project) persisted even after five iterations. Therefore, we limit feedback to a single iteration to efficiently address minor issues without incurring excessive overhead.
3.5.2 Exploiter
In this section, we systematically select the most optimal configuration of LLMs and prompts for the Exploiter.
(a) Developer Agent
(b) Critic Agent
Figure 6: Exploiter Optimal LLM Evaluation
Candidate Selection. Generating a working exploit from a given CVE in a pre-configured vulnerable environment is a crucial task for reproducing CVEs. However, the Exploiter requires less code base exploration than the Builder, as advisories often provide PoCs, exploitation steps, or vulnerability details. Thus, we include expensive reasoning models like o3, claude-3.7-sonnet, and gemini-2.5-pro in the candidates list of developers. To mitigate instability seen in open-source models (Section 3.5.1), we employ only closed-source LLMs as developers, while using the same critic LLMs as the Builder.
Developer Evaluation. For the 32 successful setups from Section 3.5.1, we ran CVE-Genie up to the Exploit Developer stage 6 and manually evaluated each developer LLM as follows: (1) we checked whether the LLM executed a PoC or command sequence that demonstrated the exploit, (2) we verified that the vulnerability was clearly triggered (e.g., uncontrolled recursion in CWE-674 resulting in a RecursionError), (3) we assessed fidelity to the CVE description, (4) we ensured no mock-up environments or fake exploits were used, and (5) we confirmed the exploit script followed the expected format (see Section 3.3.1). Among all models, o3 consistently performed best, successfully reproducing the highest number of exploits (13 out of 27) and demonstrating them end-to-end within the actual vulnerable environment. In contrast, o4-mini often produced incorrect or mock-up exploits and struggled with runtime errors. While claude and gemini models showed strong capabilities in demonstrating exploits, they lacked versatility across different vulnerability types. Hence, we select o3 as the Exploit Developer.
Critic Prompt Design. We examined the failed exploit attempts and categorized the recurring error patterns, which directly mirrored our manual verification criteria (e.g., lack of exploit demonstration, unclear vulnerability trigger, deviation from CVE description, reliance on dummy/fake exploits, or incorrect script formatting). We then incorporated these categories into the critic agent’s prompt, enabling it to automatically detect these issues for Exploit Developer.
Critic Evaluation. As shown in Figure 6(b), o4-mini achieved the best balance of recall and specificity. In contrast, o3 was overly strict, rejecting correct exploits and focusing too much on code formatting, while gemini-2.5-pro frequently overlooked incomplete exploit verification. Claude-3.7-sonnet was too lenient, accepting nearly all exploits. Based on this evaluation, we selected o4-mini as the most effective critic model for the Exploit Critic.
Feedback Assessment. We evaluated feedback effectiveness by running CVE-Genie on 32 correct setups with the top developer (o3) and critic (o4-mini). This yielded valid exploits for 16 CVEs, of which 3 were fixed based on the critic’s feedback. As in Section 3.5.1, the developer mainly resolved minor issues (e.g., incomplete verification or missing log triggers), while more complex flaws (e.g., incorrect exploit verification) persisted. To balance usefulness and cost, we therefore restrict feedback to a single iteration.
3.5.3 Verifier
In this section, we systematically select the most optimal configuration of LLMs and prompts for the CTF Verifier.
(a) Developer Agent
(b) Critic Agent
Figure 7: Verifier Optimal LLM Evaluation
Candidate Selection. Like exploit generation, creating a verifier also demands deep understanding of security vulnerabilities to ensure the exploit is successfully triggered. Therefore, we use the same developer and critic candidates as in the Exploiter.
Developer Evaluation. For the 16 valid exploits from Section 3.5.2, we ran CVE-Genie up to the Flag Checker stage 9 and manually evaluated each developer LLM on three criteria: (1) whether the verifier script followed the required format (see Section 3.4.1), (2) whether it executed the provided exploit script without modification, and (3) whether its verification logic was precise and reliable (see Figure 4 for a detailed example).
As shown in Figure 7(a), the o3 model generated 13 valid verifiers out of 14, outperforming all others. While o4-mini achieved comparable numbers, its verifiers more often relied on weak logic. Other models struggled with functional correctness and error recovery. We therefore selected o3 for the Verifier Developer due to its consistent one-shot success in producing correct verifiers.
Critic Prompt Design. We analyzed failures in verifier generation and found that the recurring error types matched our verification criteria (e.g., incomplete checks, incorrect validation logic, or misuse of the target environment). We used these categories to design the critic agent’s prompt, enabling it to automatically flag such errors when evaluating verifiers produced by the Verifier Developer.
Critic Evaluation. As shown in Figure 7(b), only the o3 model successfully identified critical but subtle issues in the verifier scripts, such as weak or flawed verification logic (Figure 4). Other models provided weak or no critique, making o3 the most
Feedback Assessment. We evaluated CVE-Genie on 16 correct exploits from the Exploiter, successfully generating accurate verifiers for 15 of them, of which 4 verifiers were improved through the critic’s feedback. In our experiment, the maximum number of feedback iterations required to resolve either functional or critique-based issues was 4; therefore, we allow up to 5 retries for both the Flag Checker and the Verifier Critic.
3.5.4 Time and Cost Constraints
In our evaluation of CVE-Genie, we found that successfully reproduced CVEs typically required about \$2 and 18 minutes on average, with the most resource-intensive case costing \$6 and taking 45 minutes. Failed reproductions, on the other hand, averaged up to \$4 and 35 minutes. To balance flexibility with cost control, we set a per-CVE budget cap of \$5 and a maximum runtime of 45 minutes for all experiments.
4 Experimental Methodology
In this section, we conduct a four-part evaluation of CVE-Genie: (1) a baseline study assessing consistency and efficiency (Section 4.1); (2) an ablation study to assess the impact of key components of CVE-Genie on its scalability and effectiveness (Section 4.2); (3) robustness testing under incomplete CVE information (Section 4.3); and (4) large-scale evaluation on 841 CVEs (Section 4.4). These evaluations are guided by the following research questions:
  • RQ1: Does the performance of CVE-Genie vary over multiple runs of CVE reproduction?
  • RQ2: Is the complex architecture of CVE-Genie necessary, or can standalone LLMs achieve comparable results?
  • RQ3: Can CVE-Genie effectively reproduce CVEs with limited information, and which CVE report components are most crucial for reproduction?
  • RQ4: Can CVE-Genie reproduce a broad range of CVEs across diverse CWEs, languages, and projects?
4.1 Baseline Assessment
In this section, we evaluate the baseline performance of CVE-Genie and examine how its reproduction success evolves across multiple runs in the presence of LLM non-determinism.
Dataset. For this study we use the same $D_{S}$ dataset (Section 3.5).
Methodology. Due to the non-deterministic nature of LLMs, a single run may fail due to incorrect reasoning, while later attempts might succeed [33]. To account for this, we run CVE-Genie multiple times on $D_{S}$, using its most optimal configuration (Section 3.5). After each iteration, we store successfully reproduced CVEs and rerun on the remaining ones. This iterative approach helps maximize CVE reproduction and reveals CVE-Genie’s convergence.
Results. We perform manual analysis of randomly selected 25 CVE reproduction runs, and observe that the variability in CVE-Genie’s performance primarily stems from the Builder phase, which is more open-ended than the targeted exploit generation guided by CVE context. Project setup involves repository exploration (Pre-Requisite Developer) and command execution (Setup Developer), where environment-specific behaviors often cause failures. For example, one run failed because the Setup Developer modified the PATH variable, while a clean retry succeeded, highlighting the value of reattempting in a fresh environment. Since the CTF Verifier depends on the Exploiter, which in turn relies on the Builder, this initial project setup is a critical bottleneck. As a previously failed build, if later successful, can enable downstream modules to reproduce a new CVE (as illustrated in Figure 8). Over multiple runs, the number of successful project, exploit, and verifier builds showed consistent gains with convergence at run seven, indicating that most recoverable failures can be resolved within a few retries.
RQ1: Due to the non-determinism of LLMs, CVE-Genie exhibits variability across runs, primarily during project setup and environment configuration. Once project setup succeeds, downstream exploit generation and verification are more stable and converge in a few retries.
Figure 8: CVE-Genie performance over five runs on $D_{S}$.
4.2 Importance of CVE-Genie’s Architecture
Having evaluated the selection of LLMs and prompts for CVE-Genie (Section 3.5), we now assess CVE-Genie’s architecture via a systematic ablation study that isolates the impact of CVE-Genie’s individual components on end-to-end CVE reproduction.
Dataset. We perform this ablation study using 15 CVEs from $D_{S}$, each with complete information (i.e., CVE description, PoC, patch commit, and security advisory) and reproduced consistently in the first three iterations of baseline assessment (Section 4.1), and call this dataset $D_{R}$. By keeping the CVE context complete, we ensure that any performance differences arise solely from modifications
Methodology. As Data Processor, Setup Developer, Exploit Developer, and Verifier Developer are the four indispensable components of CVE-Genie for end-to-end CVE reproduction, for this evaluation of CVE-Genie’s design we use the following five ablation settings:
  1. 1.
    No Knowledge Builder: Bypass the Knowledge Builder and feed raw data directly to all agents.
  2. 2.
    No Pre-Requisite Developer: Eliminate the Pre-Requisite Developer, allowing the Setup Developer to independently plan and execute the entire environment setup from scratch.
  3. 3.
    No Feedback Loops: Enforce single-shot execution without iterative refinement, by removing all feedback loops.
  4. 4.
    No Critics: Remove the Verifier Critic and treat a CVE as reproduced if the final scripts pass the Flag Checker.
  5. 5.
    Single Monolithic Agent: Combine all modular agents into a single agent with access to the same tools, thereby evaluating the performance of a standalone LLM without CVE-Genie’s agentic design and structured guidance (see Appendix G).
Results. In our ablation study (Table 4), removing the Knowledge Builder component reduces reproduction success to 9/15, as agents struggle to interpret unstructured advisory data and hit context window limits more often, which results in a 30% increase in exploit failures. Eliminating the Pre-Requisite Developer has a comparatively mild effect (13/15), but places additional burden on the Setup Developer as it has to do both codebase exploration and project setup at the same time, leading to 27% more tool-call limit violations. We observe this behavior particularly in complex repositories such as CVE-2025-32389 in Nameless, a website software for Minecraft servers, with not very clear setup guides. Feedback loops emerged as the most critical design element, and without them the reproduction success dropped sharply to 5/15 (–67%). Similarly, removing critic agents reduced success to 8/15 while increasing false reproductions by 47%, highlighting their role in reliable end-to-end CVE reproduction. Finally, collapsing all components into a single monolithic agent resulted in zero successful reproductions, even when using advanced standalone LLMs (o3), underscoring the necessity of CVE-Genie’s modular, feedback-driven, and critic-guided
RQ2: Standalone LLMs are currently unable to reproduce CVEs end-to-end. CVE-Genie’s modular, feedback-driven, and critic-guided architecture design is essential for reliable CVE reproduction, and removing any of its components significantly degrades performance (Table 4).
Ideal Case Knw Build Pre- Reqs Feed-back Critic Agents Single Agent
15 / 15 9 / 15 13 / 15 5 / 15 8 / 15 0 / 15
- $\uparrow$ 30% Exploit Failure $\uparrow$ 27% Max Tool $\downarrow$ 67% Reprod Rate $\uparrow$ 47% False Reprod -
Table 4: Ablation study of CVE-Genie’s design.
4.3 Robustness to Loss of CVE Context
To evaluate how CVE-Genie performs when critical components of a CVE report are missing, we study its robustness under systematically reduced CVE context. We first present a detailed qualitative case study of a single CVE to illustrate how CVE-Genie adapts its reasoning under different context losses, and then conduct a quantitative ablation study to quantify the impact at scale.
4.3.1 Case Study for CVE-2024-4340.
We begin by performing a detailed case study for CVE-2024-4340 reproduction under three levels of CVE context losses, and compare their reproductions against the ideal setting shown in Figure 2.
1) No Security Advisory / PoC. In this setting, all security advisories are removed from the CVE context, leaving only the developer patch commit [^2] CVE-2024-4340 Developer Patch: https://github.com/andialbrecht/sqlparse/commit/b4a39d9850969b4e1d6940d32094ee0b42a2cf03 and the CVE description [^3] CVE-2024-4340 CVE Description: “Passing a heavily nested list to sqlparse.parse() leads to a Denial of Service due to RecursionError.”. As a result, CVE-Genie relies heavily on analyzing the patch commit, exploring modified files, and inferring the PoC from the regression test added by the developer in tests/test_regressions.py. This test closely resembles the PoC described in the existing advisory [^4] CVE-2024-4340 GHSA: https://github.com/advisories/GHSA-2m57-hf25-phgg, and the generated exploit payload (Figure 9(a)) is comparable to the ideal run shown in Figure 3(a). Due to the absence of explicit PoC information, this run required additional exploration, resulting in a longer execution time (9 minutes) and higher cost (\$0.67) compared to the ideal scenario.
2) No Patch. Here, the developer patch commit is removed while security advisories remain available. CVE-Genie primarily leverages the PoC provided in the advisory, which includes clear PoC exploit code, execution traces, vulnerable code, and vulnerability root cause. Consequently, the reproduction closely matches the ideal run in terms of generated PoC exploit (Figure 9(a)), verification scripts, execution time, and cost.
3) Only CVE Description. In the most constrained setting, CVE-Genie has access only to the CVE description. Based on this limited information, it starts by identifying the implementation of sqlparse.parse and hypothesizes on how a deeply nested inputs can trigger a recursive execution path leading to a RecursionError. During PoC exploit generation and verification, the first run failed to produce a payload that successfully triggered the RecursionError. However, a second run from scratch succeeded in generating a valid payload (shown in Figure 9(b)). This setting proved to be the most challenging, requiring 20 minutes of execution time and \$4.50 LLM cost over two runs.
1payload = ’[’ * depth + ’]’ * depth
2sqlparse.parse(payload)
(a) Security advisory/PoC or patch commit is not available.
1payload = ’select’ + ’(’ * depth + ’1’ + ’)’ * depth
2sqlparse.parse(payload)
(b) Only CVE description is available.
Figure 9: PoC exploit payload generated by CVE-Genie for CVE-2024-4340 under varying levels of incomplete CVE context compared to the payload in ideal scenario in Figure 3(a).
4.3.2 Evaluation at Scale.
In this section, we perform an ablation study to systematically quantify the robustness of CVE-Genie to incomplete CVE context at scale.
Dataset ($D_{R}$). We perform this ablation with $D_{R}$ dataset (Section 4.2). We intentionally select these “easier” CVEs as a controlled probe, since CVE-Genie can consistently reproduce them under full context. Hence, any performance degradation after removing specific information indicates that the missing context is genuinely critical. To simulate incomplete CVE context, we construct three variants of $D_{R}$ by systematically removing: (a) security advisory/PoC, (b) patch commit, and (c) everything except CVE description.
Methodology. In this study, for each variant of $D_{R}$, we run CVE-Genie three times per CVE, iteratively.
Results. Removing the developer patch commit leads to a notable drop in reproduction success 10/15 (67%), even when a PoC is available. Patch commits provide essential localization signals by identifying relevant files and code paths. Without patch commits, the model must rely on broad codebase exploration, often resulting in inefficient search and budget exhaustion. We observed this behavior in CVE-2025-31481 reproduction, an authorization flaw in the PHP-based API Platform Core, where the absence of a patch caused CVE-Genie to explore unrelated components and fail to converge. Excluding security advisories reduces the success rate to 11/15 (73%). In these cases, failures are primarily attributable to environment setup and verification difficulties rather than exploit generation. Advisories often provide crucial contextual information, such as dependency versions, execution constraints, and validation guidance, that streamline reproduction, particularly when PoCs are not provided as executable code but rather as high-level textual descriptions. Despite these challenges, CVE-Genie maintains baseline robustness under minimal context. When only the CVE description is provided, it successfully reproduces 9/15 cases (60%), albeit with increased cost and runtime. This demonstrates that while comprehensive CVE reports significantly improve efficiency and reliability, meaningful reproduction remains feasible even
RQ3: CVE-Genie can reproduce CVEs under limited context. Code-based PoCs provide the strongest guidance, but in their absence, patch commits help localize vulnerabilities. Reproduction remains possible even with only CVE descriptions, but with lower success rates, higher cost, and increased time.
Figure 10: Breakdown of CVE-Genie’s run on $D_{L}$
Project Type CVEs Reprod. Projects Reprod.
AI/ML Platform 22/38 (57.9%) 14/16 (87.5%)
Web Application/Backend 171/383 (44.6%) 86/156 (55.1%)
Library/Framework 163/224 (72.8%) 115/147 (78.2%)
Cloud/DevOps Application 13/33 (39.4%) 13/25 (52.0%)
Operating System/Runtime 15/31 (48.4%) 5/15 (33.3%)
Desktop Application 6/34 (17.6%) 2/13 (15.4%)
CLI Tool/Utility 18/37 (48.6%) 15/28 (53.6%)
Blockchain/Crypto/FinTech 1/13 (7.7%) 1/7 (14.3%)
Mobile Application/SDK 2/7 (28.6%) 2/7 (28.6%)
Embedded/Networking 11/27 (40.7%) 10/18 (55.6%)
Security Tool/Server 6/12 (50.0%) 4/6 (66.7%)
Table 5: CVEs reproduction coverage by project type.
CWE-ID Overall success PoC available No PoC
CWE-79 65/142 (45.8%) 43/94 (45.7%) 22/48 (45.8%)
CWE-200 20/43 (46.5%) 15/20 (75.0%) 5/23 (21.7%)
CWE-22 16/37 (43.2%) 11/22 (50.0%) 5/15 (33.3%)
CWE-284 19/36 (52.8%) 16/26 (61.5%) 3/10 (30.0%)
CWE-400 22/36 (61.1%) 13/19 (68.4%) 9/17 (52.9%)
CWE-20 13/29 (44.8%) 6/15 (40.0%) 7/14 (50.0%)
CWE-770 14/24 (58.3%) 7/9 (77.8%) 7/15 (46.7%)
CWE-918 7/22 (31.8%) 6/13 (46.2%) 1/9 (11.1%)
CWE-89 11/21 (52.4%) 8/15 (53.3%) 3/6 (50.0%)
CWE-1333 15/19 (78.9%) 8/10 (80.0%) 7/9 (77.8%)
CWE-94 9/19 (47.4%) 2/8 (25.0%) 7/11 (63.6%)
CWE-863 7/15 (46.7%) 4/6 (66.7%) 3/9 (33.3%)
CWE-287 8/15 (53.3%) 4/4 (100.0%) 4/11 (36.4%)
CWE-285 7/15 (46.7%) 5/8 (62.5%) 2/7 (28.6%)
CWE-269 7/13 (53.8%) 4/8 (50.0%) 3/5 (60.0%)
CWE-122 8/13 (61.5%) 2/4 (50.0%) 6/9 (66.7%)
CWE-502 6/13 (46.2%) 3/6 (50.0%) 3/7 (42.9%)
CWE-532 6/12 (50.0%) 2/2 (100.0%) 4/10 (40.0%)
CWE-78 4/11 (36.4%) 4/9 (44.4%) 0/2 (0.0%)
CWE-352 8/10 (80.0%) 5/6 (83.3%) 3/4 (75.0%)
CWE-74 7/10 (70.0%) 4/4 (100.0%) 3/6 (50.0%)
CWE-416 3/10 (30.0%) 0/0 (–) 3/10 (30.0%)
CWE-639 2/9 (22.2%) 1/4 (25.0%) 1/5 (20.0%)
CWE-116 4/8 (50.0%) 1/3 (33.3%) 3/5 (60.0%)
CWE-347 7/8 (87.5%) 4/4 (100.0%) 3/4 (75.0%)
Table 6: Top 25 CWEs identified in the large-scale study and their reproduction success rates, overall and stratified by PoC availability.
Project # Re CVEs # T25 CWEs # Oth CWEs LoC Lang
lunary-ai 35 13 22 ~70k TypeScript
WeGIA 17 15 2 ~1M PHP
cpython 10 2 8 ~1.8M Py, C
vite 6 6 0 ~107k TypeScript
yeswiki 6 6 0 ~234k PHP
symfony 5 4 1 ~1.6M PHP
llama_index 5 4 1 ~1.1M Py
vim 5 3 2 ~1.2M Vim Script, C
directus 5 2 3 ~398k TypeScript
rails 4 2 2 ~450k Ruby
phpipam 4 2 2 ~225k PHP
assimp 4 4 0 ~562k C++
siyuan 3 1 2 ~265k Go, TypeScript
cvat-ai 3 3 0 ~320k Py, TypeScript
MobSF 3 1 2 ~132k JS, Py
mlflow 3 1 2 ~870k Py
Table 7: Top 15 projects identified in the large-scale evaluation of CVE-Genie, showing the number of reproduced CVEs, their classification into CWE Top 25 vs. other CWEs, lines of code (LoC), and primary languages.
4.4 Large Scale Evaluation
Dataset $(D_{L})$. For our large-scale evaluation, we construct a dataset, $(D_{L})$, consisting of all CVEs published between June 2024 and May 2025 that include publicly available source code. A CVE is included in $(D_{L})$ if at least one of the following conditions is satisfied: (1) the CVE entry contains a GitHub URL (e.g., a patch commit) in its references that links directly to the relevant source code, or (2) the CVE specifies affected software versions, and the source code corresponding to those versions is publicly accessible. Because these CVEs were disclosed after the knowledge cutoff of the LLMs (o3 and o4-mini) used in CVE-Genie, $(D_{L})$ ensures zero prior exposure during model training. Overall, $(D_{L})$ comprises 841 CVEs spanning 186 CWEs, 29 programming languages, and 440 distinct open-source software (OSS) projects. The OSS projects cover a broad range of application domains, including web applications and backends (156), libraries and frameworks (147), CLI tools and binaries (28), cloud and DevOps systems (25), embedded and networking software (18), operating systems and runtimes (17), AI/ML platforms (16), desktop applications (13), blockchain and cryptocurrency systems (7), mobile applications and SDKs (7), and security tools (6). In terms of vulnerability types, the majority of CVEs in $(D_{L})$ fall under the following categories: injection and improper neutralization (29%), authorization and authentication (11%), resource management (8%), information exposure (6%), memory safety (3%), request forgery (3%), and others. Our statistical analysis further shows that approximately 55% of CVEs in $D_{L}$ fall under the top 25 most dangerous CWEs [37], and around 49% lack a PoC in their advisory.
Methodology. We use the same methodology as Section 4.1, and run CVE-Genie on $D_{L}$ three times, iteratively.
Results. CVE-Genie reproduced 428 CVEs out of 841 (Figure 10), spanning 22 programming languages, 141 vulnerability types, and 267 projects. To better understand outcomes, we perform statistical analysis on all reproduction runs as well as manually analyze 50 successful and 50 failed cases, and categorize results as follows:
Domain-Dependent Effectiveness of CVE-Genie. We observe that CVE-Genie shows the highest performance for library frameworks and AI/ML platforms categories, with the reproduction rates of 78% and 58%, excelling on CWEs such as CWE-79 (XSS), CWE-89 (SQLi), and CWE-284 (access control). Meanwhile, the success appears more moderate for web/backend applications (45%), where logic-driven CWEs such as 200, 22, and 918 have lower success rates. CVE-Genie struggles in binary-heavy domains such as desktop (18%), blockchain (8%), and mobile (29%) applications, especially with memory-safety (e.g., CWE-122, 416) and crypto/validation (CWE-295, 327) flaws. A finer-grained breakdown by vulnerability type and project characteristics is provided in the Appendix, where Table 6 reports reproduction success rates for the top CWEs. Meanwhile, Table 7 and Table 5 further contextualize results by project scale, language, and domain. Finally, Tables 919 detail domain-specific CWE distributions, illustrating how failure
Failures. For $198$ cost/time overrun failures, ~80% were due to cost overruns, where the agent either got stuck during project builds (~41%) or while generating exploits (~59%). The majority of failures involved web application CVEs lacking either vulnerability details or setup documentation. For example, none of the CVEs in project discourse built successfully since its README only links to external documentation, which our agents cannot access, causing the exhaustion of their budgets in exploring the codebase.
RQ4: At scale, CVE-Genie reproduced 428 / 841 CVEs, spanning 267 projects, 22 programming languages, and 141 CWEs, demonstrating broad applicability across ecosystems. Notably, 217 (51%) reproduced CVEs fall within the MITRE top 25 most dangerous CWEs, and 231 (54%) had no PoC available.
5 Limitations and Future Work
Vulnerabilities Involve UI Interactions. Currently, CVE-Genie only supports command-line interface (CLI) interactions. However, many web-based vulnerabilities require interaction with a graphical UI, making them difficult to trigger via CLI. Future work should explore integrating CVE-Genie with UI environments, e.g., web browsers.
Multimodal CVE Knowledge Curation. Due to a lack of image and video processing support, CVE-Genie struggles with CVEs where the PoC is provided as an image or video. Future work can address this limitation by integrating CVE-Genie with multimodal LLMs. Additionally, leveraging external CVE resources through LLM-based deep search tools, beyond standard vulnerability databases, may also help improve reproduction success rates [38].
Over Critique. As shown in Figures 5(b), 6(b), and 7(b), all critics exhibit a lower TPR compared to their TNR. While this helps reduce false positives, it also results in missing some valid reproductions, particularly in cases where critics requested detailed setup, exploit, or verifier that the developer could not sufficiently provide.
Cost. Running CVE-Genie costs approximately \$2.77 per CVE, moreover, open-source models currently do not perform well. Therefore, to support further research, we release logs from all reproduction runs as a novel dataset to help the community fine-tune and improve local models’ performance for CVE reproduction.
6 Conclusion
In this paper we presented CVE-Genie, an automated framework designed to reproduce real-world vulnerabilities at scale. This has considerable benefit for downstream security research (see Appendix F), which suffers from a shortage of diverse high-quality data suitable for research and development of new tooling for automated vulnerability discovery and repair. CVE-Genie successfully reproduced around 51% (428) of 841 CVEs, with on average \$2.77 cost per CVE, across a vast variety of projects, CWEs, and programming languages. To the best of the authors’ knowledge, no other framework is capable of such comprehensive results at this scale and efficiency.
Appendix A Acknowledgments
We gratefully acknowledge and thank Dave Aitel, Harold Nguyen, Xin Hu, and Ian Brelinsky (OpenAI), as well as Stijn Pletinckx and Lukas Dresel (University of California, Santa Barbara), for their valuable discussions and guidance that contributed to this work.
Appendix B Ethical Considerations
Safe Release of Reproduced CVEs’ Data. For all of our studies, we only include the CVEs that have a public patch available on the National Vulnerability Database (NVD). This measure prevents the attackers from using our reproduced CVEs to exploit the stakeholders’ systems, and also helps us to safely release the data of the CVEs reproduced in our experimental studies.
Release of CVE-Genie’s Codebase. The aim of releasing CVE-Genie’s codebase is to support defenders and solution developers by accelerating vulnerability triaging and tool development for vulnerability detection and patching. Although it is also possible for the attackers to leverage CVE-Genie to reproduce vulnerabilities in the real world, our aim is to level the playing field for defenders; specifically, to empower defenders to reproduce CVE reports to attain up-to-date and reproducible CVE environments as benchmark test cases. These benchmarks will then be used to rigorously test and enhance vulnerability detection and patching systems, thereby understanding their performance and ideally accelerating their response capabilities. Following the beneficence criterion of the Menlo Report [14], we strongly believe that our system provides more benefits than potential harms to the security community and to society.
Appendix C Open Science
CVE-Genie Framework and Dataset. https://github.com/BUseclab/cve-genie. This includes (1) source code of CVE-Genie and detailed instructions on how to run it, (2) reproduction runs of all reproduced CVEs with their agent conversations/logs, intermediate artifacts, and final exploit and verifier scripts, and (3) a web application to visualize the CVE reproduction runs.
Experiments Results. https://osf.io/dcej4/overview?view_only=81309bbc7e8c489abb9fb7e40bd1c0fb. This includes all reproduction runs with their agent conversations/logs, intermediate artifacts, and final scripts. (1) critics_evals.tar.gz contains results for selecting optimal LLMs and prompts (Section 3.5). (2) baseline.tar.gz contains results for the baseline assessment study (Section 4.1). (3) design_ablation.tar.gz contains results for CVE-Genie’s design ablations (Section 4.2). (4) robustness.tar.gz contains results for CVE-Genie’s robustness to loss of CVE context (Section 4.3). (5) large_scale.tar.gz contains results for multiple runs of the large-scale study with 841 CVEs (Section 4.4).
Appendix D Generative AI Usage
We used ChatGPT (OpenAI) to assist with LaTeX formatting, as well as to review and refine spelling, grammar, and sentence structure for improved clarity of the manuscript. Moreover, all AI-assisted outputs were manually reviewed, edited as needed, and explicitly approved by the authors before being incorporated into the paper.
Appendix E CVE-Genie Specifications
CVE-Genie is built on LangChain and supports easy design and integration of tools for LLMs. For CVE-Genie to effectively reproduce CVEs, it require access to a project’s source code. However, due to LLMs’ limited context windows, large project files cannot be processed all at once. To address this, we developed primitives enabling directory browsing using “project directory tree” and command execution:
  1. 1.
    get_file: The LLM agent specifies the file path and the number of lines it wants to read from a given offset. To prevent context overflow and improve efficiency, it reads a maximum of 300 lines at a time. If more content is needed, the agent can scroll up or down to access it.
  2. 2.
    write_to_file: The LLM provides the path and content for a file, which we then write to the specified file. This facilitates modifications and updates within the project.
  3. 3.
    execute_ls_command: The LLM specifies a directory, and we execute the ls command within it, returning the output. This command is distinct from execute_linux_command due to its major role in project directory exploration and its frequent use.
  4. 4.
    execute_linux_command: The LLM issues Linux commands to be executed from the project’s root directory. Foreground commands have a 300-second timeout, with standard output and error logged to separate files. After completion, we return up to the last 100 lines of the log and the log file path which LLM can read using get_file for further analysis. Background execution is supported for non-blocking tasks (e.g., starting a web server). For these, we show output after 5 seconds, provide the log file path, and indicate whether the process is still running.
  5. 5.
    set_environment_variable: This tool allows for setting environment variables for subsequent execute_linux_command calls. We observed that LLMs frequently issue calls to export environment variables, which are not retained in successive runs as each command is executed in a separate shell. To address this, we enable the LLM to set these variables, which are then passed to commands. There is also an option to clear all set variables.
Restrictions. Each tool-calling agent is limited to 60 tool calls. This limit was found to be optimal in preliminary experiments for maximizing CVE reproduction across all LLMs.
Output Formatting. For all agents requiring output in a specific format, we add a gpt-4o-mini-based format corrector to the output parser. If parsing fails, the raw output is passed to gpt-4o-mini for formatting. If formatting fails after three attempts, an error is returned.
Category Application Benefits brought by CVE-Genie
Vul. detection ML-based detection Training and testing data
Dynamic fuzzing Benchmarks and initial fuzzing seeds
Static analysis Benchmarks and static rules
Vul. triage Root cause analysis Enrich inputs and enable more tools
Exploit generation Ingredients for exploit chaining
Vul. Patching Patch generation Benchmarks and testing environments
Patch verification Inputs and environments
Other applications LLM secure code generation Benchmarks and testing environment
Penetration testing Construct pen. test tasks and environment
Attack detection Benchmarks and detection rules
Table 8: Tasks that benefit from CVE-Genie.
Appendix F Applications of CVE-Genie
As shown in Table 8, reproducible CVEs can be used for a plethora of security analysis tasks, including vulnerability detection, triage, exploitation, and patching.
Vulnerability Detection. Most existing vulnerability detection datasets are built using patch commits from CVEs (see Table 1). They typically label the pre-patch version of a function as vulnerable and the post-patch version as benign [40, 15, 7]. However, extracting isolated functions often strips away critical context [51], and functions labeled for a single vulnerability may contain others, resulting in noisy labels. Using reproducible CVEs with working PoCs can address these issues. By executing PoCs during function code extraction, we can ensure that vulnerable execution traces are included, preserving necessary context and confirming the presence of vulnerabilities. Additionally, this approach supports both static and dynamic analysis, unlike most existing datasets (e.g., SVEN [21], BigVul [15], PrimeVul [13]), which are limited to static methods.
Vulnerability Triage and Patching. Reproducible CVEs are essential for effective vulnerability triage. They help eliminate false positives and provide an executable environment with known vulnerable inputs, which aids in accurately analyzing and reproducing bugs. This capability enables advanced techniques like reverse execution and backward taint analysis for root cause identification. CVE-Genie supports this process by offering exploits for individual vulnerabilities, which can be combined into complex exploit chains for studying multi-stage attacks. It also serves as a benchmark for evaluating patching methods, providing rich contextual data such as crash stack traces and outputs. Additionally, its PoCs assist in verifying and refining patches.
LLM Insecure Code Generation. Reproducible CVEs are valuable for evaluating the security of code generated by LLMs. Recent studies have shown that LLMs can produce insecure code, particularly in security-critical contexts [59, 55]. Various benchmarks evaluate secure code generation capability of LLM, such as, CyberSecEval [55] prompts an LLM to generate code snippets and uses rule-based detectors to identify insecure code, however, rule-based detection often leads to false positives. To improve this, SecCodePLT [59] leverages CWE data and manually crafts dynamic test cases to evaluate secure code generation. Reproducible CVEs offer a scalable alternative because these CVEs are confirmed to be vulnerable, they can be used to assess LLMs’ ability to generate secure code. By prompting the LLM to implement the functionalities in the vulnerable functions, and using the associated PoCs to test the output, we can determine whether the generated code is vulnerable.
Penetration Testing and Attack Detection. CVE-Genie offers a valuable resource for both penetration testing and cybersecurity training. It provides ready-to-use exploitation scenarios along with verified solutions, making it ideal for educating future penetration testers and developing automated testing tools. For defenders, these exploits can be analyzed to identify attack patterns, which can enhance detection mechanisms such as intrusion detection systems (IDS) or malware classifiers.
Appendix G Single Monolithic Agent Study Design
To evaluate a standalone LLM without CVE-Genie’s modular, agentic design and structured guidance, we collapsed all agents into a single monolithic agent with access to the full toolset (Appendix A), its entire conversation history, and an increased limit of 100 tool calls per attempt. The agent was provided with all CVE-related information directly in the user prompt, and the system prompt explicitly instructed it to follow a three-stage workflow (project setup, exploit generation, and verification), mirroring CVE-Genie’s high-level procedure. Unlike CVE-Genie, no automated critic or feedback loops were used; instead, we manually inspected execution traces and the generated exploit and verifier artifacts to determine whether reproduction was successful.
Appendix H Case Study for CVE-2024-5129
Overview. Versions of the lunary project prior to v1.2.8 expose a DELETE /v1/datasets/:id endpoint that lacks authorization checks and performs an unscoped deletion by dataset UUID. In the vulnerable release v1.2.7, the handler executes a SQL DELETE using only id, with no authentication and no restriction to the caller’s project, allowing unauthenticated deletion of arbitrary datasets by guessing or obtaining UUIDs.
CWE-IDs # CVEs PoC Available No PoC
Reprod Failed Reprod Failed
CWE-502 6 3 1 1 1
CWE-29 3 2 1 0 0
CWE-20 3 1 2 0 0
CWE-77 3 0 2 1 0
CWE-94 2 0 1 0 1
Table 9: Top CWEs reproduced in AI/ML Platform projects.
CWE-IDs # CVEs PoC Available No PoC
Reprod Failed Reprod Failed
CWE-79 99 34 37 10 18
CWE-284 29 12 8 2 7
CWE-22 21 7 6 3 5
CWE-200 19 5 3 2 9
CWE-918 16 4 5 1 6
CWE-89 15 5 6 1 3
CWE-400 15 6 3 1 5
CWE-20 12 4 3 1 4
CWE-285 9 5 2 2 0
Table 10: Top CWEs reproduced in Web App projects.
CWE-IDs # CVEs PoC Available No PoC
Reprod Failed Reprod Failed
CWE-79 33 9 6 12 6
CWE-1333 15 8 1 4 2
CWE-400 13 5 1 6 1
CWE-200 13 9 0 1 3
CWE-770 13 5 0 5 3
CWE-22 8 3 1 2 2
Table 11: Top CWEs reproduced in Lib/Framework projects.
CWE-IDs # CVEs PoC Available No PoC
Reprod Failed Reprod Failed
CWE-285 4 0 1 0 3
CWE-209 3 0 2 0 1
CWE-200 3 0 0 1 2
CWE-269 2 1 1 0 0
CWE-287 2 0 0 0 2
Table 12: Top CWEs reproduced in Cloud/DevOps projects.
CWE-IDs # CVEs PoC Available No PoC
Reprod Failed Reprod Failed
No-ID 7 0 0 5 2
CWE-122 3 0 0 3 0
CWE-400 2 0 0 1 1
CWE-416 2 0 0 0 2
CWE-863 2 0 0 0 2
Table 13: Top CWEs reproduced in OS/Runtime projects.
CWE-IDs # CVEs PoC Available No PoC
Reprod Failed Reprod Failed
CWE-79 7 0 6 0 1
CWE-122 4 0 0 1 3
CWE-416 4 0 0 2 2
CWE-22 2 0 2 0 0
CWE-20 2 0 1 0 1
Table 14: Top CWEs reproduced in Desktop App projects.
CWE-IDs # CVEs PoC Available No PoC
Reprod Failed Reprod Failed
No-ID 4 0 0 1 3
CWE-345 3 2 0 1 0
CWE-116 2 0 0 1 1
CWE-147 2 0 0 1 1
CWE-150 2 0 0 1 1
Table 15: Top CWEs reproduced in CLI Tool/Utility projects.
CWE-IDs # CVEs PoC Available No PoC
Reprod Failed Reprod Failed
CWE-670 3 1 1 0 1
CWE-200 2 0 2 0 0
CWE-400 1 0 1 0 0
Table 16: Top CWEs reproduced in Blockch/Crypto projects.
CWE-IDs # CVEs PoC Available No PoC
Reprod Failed Reprod Failed
CWE-532 2 0 0 1 1
CWE-261 1 0 0 1 0
CWE-327 1 0 0 1 0
Table 17: Top CWEs reproduced in MobileApp/SDK projects.
CWE-IDs # CVEs PoC Available No PoC
Reprod Failed Reprod Failed
CWE-200 4 1 0 1 2
CWE-416 3 0 0 1 2
CWE-670 2 1 0 0 1
CWE-639 2 1 0 0 1
Table 18: Top CWEs reproduced in Emb/Network projects.
CWE-IDs # CVEs PoC Available No PoC
Reprod Failed Reprod Failed
CWE-918 3 0 2 0 1
CWE-863 2 0 0 2 0
CWE-23 1 1 0 0 0
Table 19: Top CWEs reproduced in Security/Server projects.
Processor. The workflow starts with Data Processor resolving CVE-2024-5129’s affected versions and automatically downloading the latest vulnerable release (v1.2.7) from the upstream repository. Data Processor also extracts the CVE data, CWE information, patch commit, and bug bounty report for the CVE. The Knowledge Builderthen constructs a structured CVE knowledge base summarizing the vulnerability as missing authorization (CWE-862) and localizing the root cause to v1/datasets/index.ts, where the vulnerable delete handler deletes by dataset ID alone. The knowledge base also records the upstream fix (adding authorization middleware and scoping deletion by projectId), which becomes the ground truth for later exploit and verifier synthesis.
Builder. Given the vulnerable source tree and knowledge base, the Builder reconstructs a runnable environment. The Pre-Requisite Developer agent inspects repository structure and reads key configuration files (e.g., root package.json, backend package.json, backend .env.example, migration runner src/migrate.ts) to infer the required service stack: a Node.js/Koa backend backed by PostgreSQL and initialized via SQL migrations under packages/db.
The Setup Developer agent then provisions a real PostgreSQL 14 service inside the VM, configures credentials, and generates a valid .env file pointing DATABASE_URL to the local database instance. All Node.js dependencies are installed via npm install, and database migrations are executed using the project’s native migration runner (npm run migrate:db), resulting in a fully initialized schema. Finally, the backend is launched using npm run dev and confirmed operational by observing successful database connectivity and a listening API server on port 3333. The Setup Critic agent analyzes the execution logs of Setup Developer and approves the build.
After environment reconstruction, the Setup Critic evaluates the entire setup. The critic confirms that (i) the vulnerable code path is present and unchanged, (ii) a real PostgreSQL-backed Lunary backend is running, (iii) no dummy services or mocked endpoints were introduced, and (iv) the vulnerable DELETE /v1/datasets/:id endpoint is reachable and responds with HTTP 200 to unauthenticated requests. Based on these checks, the critic accepts the project build as a genuine and exploitable reproduction target.
Exploiter. The Exploit Developer agent begins by performing targeted code inspection to ground the exploit in the real implementation. It reads the vulnerable dataset router source file and confirms that the datasets.delete("/:id") handler executes an unscoped SQL deletion without any authentication or authorization middleware. It also inspects backend configuration files to verify the correct runtime entry points and dependencies. To validate exploitability in practice, the exploiter launches the backend and issues unauthenticated DELETE requests against /v1/datasets/<uuid> using both ad-hoc Python test scripts and raw HTTP calls. A randomly generated UUID is sufficient to elicit an unconditional HTTP 200 response, confirming that no permission checks or project scoping are enforced. Based on this confirmation, the exploiter generates a standalone PoC exploit.py script. The script accepts a dataset UUID (and an optional base URL), sends an unauthenticated DELETE request to the vulnerable endpoint, and prints the HTTP status code and response body. The exploit is intentionally minimal, containing no setup logic or auxiliary assumptions beyond network access to the API.
The generated exploit artifact is then evaluated by the Exploit Criticagent. The critic verifies that the exploit (i) targets the correct vulnerable endpoint, (ii) directly triggers the missing-authorization flaw described in the CVE knowledge base, (iii) does not rely on environmental shortcuts or fabricated behavior, and (iv) has been empirically validated against the running vulnerable service. After confirming these properties, the critic accepts the exploit as a correct and faithful PoC.
CTF Verifier. To automatically confirm successful exploitation of CVE-2024-5129, the Verifier Developer agent synthesizes a CTF-style verifier that executes the exploit and validates its effect against the reconstructed Lunary backend. The verifier follows a fixed the following three-phase structure.
  1. 1.
    In the pre-setup phase, it inserts a minimal but valid set of rows, i.e., organization, account, project, and dataset, directly into the PostgreSQL database used by the running backend. A fresh dataset UUID is generated and recorded, ensuring that a concrete deletion target exists prior to exploitation.
  2. 2.
    In the exploit phase, the verifier invokes the previously generated exploit.py script verbatim via subprocess.run, passing the dataset UUID and the base URL http://localhost:3333. The verifier itself performs no HTTP requests, ensuring that the exploit remains the sole trigger of the vulnerability.
  3. 3.
    In the post-exploit phase, the verifier re-queries the database to determine whether the dataset row with UUID still exists. If the row has been deleted, the verifier prints the flag 3xploit66full; otherwise, it reports failure. Database access is implemented robustly, using psycopg2 when available and falling back to the psql client if necessary.
During verification, earlier verifier iterations that relied on mocked or stubbed HTTP servers were rejected by the Verifier Critic agent. The final accepted verifier operates exclusively on the real Node.js / Koa backend and PostgreSQL instance and performs a direct state-based validation. The Flag Checker successfully executes the verifier, observes flag emission, and the Verifier Critic confirms the verifier as a correct and faithful reproduction of the CVE.
Storing Artifacts and Metadata. After the successful verification, CVE-Genie stores the VM snapshot with running Lunary server. Moreover, it stores exploit and verifier scripts, agent logs, and metadata of the reproduction run. Overall, this reproduction run of CVE-2024-5129 cost \$1.68 and 20 minutes. See artifacts here  [^5] https://github.com/BUseclab/cve-genie/tree/main/results/reproduced_cves/CVE-2024-5129 in open-source dataset.